### v1

# General effective reduction theory of integral polynomials of given non-zero discriminant and its applications

Abstract.
We give a survey on the general effective reduction theory of integral polynomials and its applications. We concentrate on results providing the finiteness for the number of ‘$\mathbb{Z}$-equivalence classes’ and ‘$GL_{2}(\mathbb{Z})$-equivalence classes’ of polynomials of given discriminant. We present the effective finiteness results of Lagrange from 1773 and Hermite from 1848, 1851 for quadratic resp. cubic polynomials. Then we formulate the general ineffective finiteness result of Birch and Merriman from 1972, the general effective finiteness theorems of Győry from 1973, obtained independently, and of Evertse and Győry from 1991, and a result of Hermite from 1857 not discussed in the literature before 2023 for a reason explained below. We briefly outline our effective proofs which depend on Győry’s effective results on unit equations, whose proofs involve Baker’s theory of logarithmic forms. Then we focus on our recent joint paper with Bhargava, Remete and Swaminathan from 2023, where Hermite’s finiteness result from 1857 involving ‘Hermite equivalence classes’ is compared with the above-mentioned modern results, using the two classical equivalences, and where it is confirmed that Hermite’s result from 1857 is much weaker than the modern results mentioned. Since 1973, the results of Győry from 1973 and Evertse and Győry from 1991 together established a general effective reduction theory of integral polynomials with given non-zero discriminant which resulted in many significant consequences and applications, including Győry’s effective finiteness theorems from the 1970’s on monogenic orders and number fields. For generalizations and further applications we refer to our monograph [J.-H. Evertse and K. Győry, Discriminant Equations in Diophantine Number Theory, Camb. New Math. Monogr 32, Cambridge University Press, 2017] and Sections 58 of the present paper. In the Appendix we discuss related topics not strictly belonging to the reduction theory of integral polynomials considered.
2020 Mathematics Subject Classification: 11C08, 11D72, 11J86
Keywords and Phrases: polynomials, discriminants, $\mathbb{Z}$-equivalence, $GL_{2}(\mathbb{Z})$-equivalence, Hermite equivalence, power integral bases, monogenic number fields
1. Introduction
We give an overview of older and recent results on the reduction theory of integral polynomials of given discriminant, and its many consequences and applications. We first recall some definitions and notation.
1.1. Preliminaries
a
Two polynomials $f,g\in\mathbb{Z}[X]$ of degree $n$ are called $\mathbb{Z}$-equivalent if $g(X)=f(X+a)$ or $g(X)=(-1)^{n}f(-X+a)$ for some $a\in\mathbb{Z}$, and $GL_{2}(\mathbb{Z})$-equivalent if $g(X)=\pm(cX+d)^{n}f\big{(}\frac{aX+b}{cX+d}\big{)}$ for some matrix $\big{(}\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big{)}\in GL_{2}(\mathbb{Z})$, i.e., $a,b,c,d\in\mathbb{Z}$ and $ad-bc=\pm 1$. Clearly, $\mathbb{Z}$-equivalence implies $GL_{2}(\mathbb{Z})$-equivalence. Polynomials that are $\mathbb{Z}$-equivalent to a monic polynomial are also monic.
The discriminant of a polynomial
$f=a_{0}X^{n}+\cdots+a_{n}=a_{0}\prod_{i=1}^{n}(X-\alpha_{i})$
is defined by
$D(f):=a_{0}^{2n-2}\prod_{1\leq i<j\leq n}(\alpha_{i}-\alpha_{j})^{2}.$
This is a homogeneous polynomial of degree $22n-$ 2 in $\mathbb{Z}[a_{0},\ldots,a_{n}]$; thus, if $f\in\mathbb{Z}[X]$ then $D(f)\in\mathbb{Z}$. As one may easily verify, polynomials that are $\mathbb{Z}$-equivalent or $GL_{2}(\mathbb{Z})$-equivalent have the same discriminant.
We define the height $H(f)$ of a polynomial $f=a_{0}X^{n}+\cdots+a_{n}\in\mathbb{Z}[X]$ by
$H(f):=\max(|a_{0}|,\ldots,|a_{n}|).$
Roughly speaking, reduction theory of polynomials entails to find, for a given polynomial $f\in\mathbb{Z}[X]$, another polynomial $g\in\mathbb{Z}[X]$ that is $GL_{2}(\mathbb{Z})$-equivalent (or $\mathbb{Z}$-equivalent in the monic case) to $f$ and whose coefficients have as small as possible absolute values. In this paper, we focus on results in which the height $H(g)$ of $g$ is bounded above in terms of $|D(f)|$, i.e., on reduction theory for polynomials of given discriminant. Such results imply that up to $GL_{2}(\mathbb{Z})$-equivalence (resp. $\mathbb{Z}$-equivalence if we restrict ourselves to monic polynomials) there are only finitely many polynomials $f\in\mathbb{Z}[X]$ of degree $n$ and given discriminant $D\not=0$.
The results on reduction theory for univariate polynomials can be translated immediately into similar results for binary forms. We decided to formulate our results in terms of univariate polynomials for convenience of presentation.
1.2. Brief summary of results
a
The reduction theory of integral polynomials of given non-zero discriminant was initiated by Lagrange (1773). For quadratic polynomials he proved that up to the classical $GL_{2}(\mathbb{Z})$-equivalence, resp. $\mathbb{Z}$-equivalence (monic case) there are only finitely many quadratic polynomials in $\mathbb{Z}[X]$ of given discriminant. Lagrange’s result is effective in the sense that one can effectively determine the polynomials.
Hermite (1848, 1851) introduced a reduction theory for polynomials of arbitrary degree but using another invariant instead of the discriminant, which for cubic polynomials gives the analogue of Lagrange’s result. Hermite was apparently interested to extend this to the general case, i.e., that for every $n\geq 4$ and $D\not=0$ there are up to $GL_{2}(\mathbb{Z})$-equivalence only finitely many polynomials $f\in\mathbb{Z}[X]$ of degree $n$ and discriminant $D$. In Hermite (1857) he introduced a new equivalence relation (called by us ‘Hermite equivalence’, see Section 3) and proved in an ineffective way a finiteness result on the corresponding equivalence classes of integral polynomials of degree $n$ and discriminant $D$. But he did not compare his equivalence relation to the classical equivalence relations, i.e., to $GL_{2}(\mathbb{Z})$-equivalence and $\mathbb{Z}$-equivalence. The result of Hermite (1857) does not appear to have been studied in the literature until the excellent book of Narkiewicz (2018), where Hermite equivalence was mixed up with the classical equivalence relations.
Hermite’s apparent goal, i.e., the finiteness result with $GL_{2}(\mathbb{Z})$-equivalence instead of Hermite equivalence, was finally achieved more than a century later by Birch and Merriman (1972) for arbitrary polynomials in an ineffective form and independently, for monic polynomials and in a more precise and effective form by Győry (1973). The result of Birch and Merriman was subsequently made effective by Evertse and Győry (1991). More precisely, Győry (1973) and Evertse and Győry (1991) proved that there exists an effectively computable number $c(D)$ depending only on $D$ such that every $f\in\mathbb{Z}[X]$ of discriminant $D\not=0$ is $GL_{2}(\mathbb{Z})$-equivalent (and even $\mathbb{Z}$-equivalent in the monic case) to a polynomial $g$ whose coefficients have absolute values bounded above by $c(D)$. These results heavily depend on effective finiteness results for unit equations $ax+by=1$ with solutions $x,y$ from the unit group of the ring of integers of a number field, which were derived in turn using Baker’s theory of logarithmic forms. This solved the old problem of Hermite (1857) mentioned above in an effective way, and further resulted in many significant consequences and applications. For example, in the 1970’s, Győry deduced from his paper from 1973 the first general effective algorithm that decides monogenicity and existence of power integral bases of number fields, and in fact finds all power integral bases. For later applications and generalizations we refer to the monograph Evertse and Győry (2017) and Sections 58 of the present paper.
In our recent paper BEGyRS (2023) with Bhargava, Remete and Swaminathan we provided a thorough treatment of the notion of Hermite equivalence, and proved that $\mathbb{Z}$-equivalence and $GL_{2}(\mathbb{Z})$-equivalence are much more precise than Hermite equivalence. This confirmed that Hermite’s result from 1857 was much weaker than those of Birch and Merriman, Győry, and that of Evertse and Győry mentioned above. It should of course be mentioned that unlike the last authors, Hermite didn’t have the powerful Baker’s theory of logarithmic forms at his disposal.
In the remainder of the paper we discuss in more detail the results of Birch and Merriman (1972), Győry (1973), Evertse and Győry (1991), and those from the paper BEGyRS (2023). In the Appendix we go into some related topics not strictly belonging to reduction theory of integral polynomials. In a later, yet to be written extended version of our present paper we shall publish Section B from the Appendix, which deals with so-called rationally monogenic orders, in a more detailed form.
2. Reduction theory of integral quadratic and cubic polynomials of given non-zero discriminant
As we mentioned, Lagrange (1773) was the first to develop a reduction theory for binary quadratic forms with integral coefficients. His theory was made more precise by Gauss (1801). For integral polynomials, their theories imply the following. Recall that the height $H(g)$ of a polynomial with integer coefficients is the maximum of the absolute values of its coefficients.
Theorem 2.1 (Lagrange, 1773; Gauss, 1801).
For any quadratic polynomial $f\in\mathbb{Z}[X]$ of discriminant $D\neq 0$, there exists $g\in\mathbb{Z}[X]$, $GL_{2}(\mathbb{Z})$-equivalent to $f$, such that $H(g)\leq c(D)$ with some effectively computable constant $c(D)$ depending only on $D$.
For monic polynomials, the following more precise variant is known.
Theorem 2.2.
For any monic quadratic polynomial $f\in\mathbb{Z}[X]$ of discriminant $D\neq 0$, there exists $g\in\mathbb{Z}[X]$, $\mathbb{Z}$-equivalent to $f$, such that $H(g)\leq c^{\prime}(D)$ with some effectively computable constant $c^{\prime}(D)$ depending only on $D$.
The above results have the following effective equivalent variants.
Theorem 2.3.
There are only finitely many $GL_{2}(\mathbb{Z})$-equivalence (resp. $\mathbb{Z}$-equivalence) classes of quadratic (resp. monic quadratic) polynomials in $\mathbb{Z}[X]$ of given discriminant $D\neq 0$. Further, each equivalence class has a representative of height at most $c(D)$ (resp. $c^{\prime}(D)$).
Later, mostly these equivalent versions were investigated, used and generalized.
Hermite (1848, 1851) studied integral binary forms of degree larger than $2$. He developed an effective reduction theory for such forms which implies, among other things, the following:
Theorem 2.4 (Hermite, 1848, 1851).
There are only finitely many $GL_{2}(\mathbb{Z})$-equivalence classes of cubic polynomials in $\mathbb{Z}[X]$ of given non-zero discriminant, and a full set of representatives of these classes can be effectively determined (in the sense that the proof provides an algorithm to determine, at least in principle, a full system of representatives).
In fact, Hermite (1848, 1851) introduced another invariant for polynomials $f\in\mathbb{Z}[X]$ of arbitrary degree, which is in fact the discriminant $\Delta_{f}$ of a positive definite binary quadratic form $\Phi_{f}(X,Y)=AX^{2}+BXY+CY^{2}\in\mathbb{R}[X]$ associated with $f$. He called $f$ reduced if $\Phi_{f}$ is reduced in Gauss’ sense, i.e., if $|B|\leq A\leq C$. He showed that $f$ is $GL_{2}(\mathbb{Z})$-equivalent to a reduced polynomial $g$, and that the coefficients of $g$ are bounded effectively in terms of $\Delta_{f}$. Hermite showed further that for cubic $f$, $\Delta_{f}=|27D(f)|^{1/4}$, implying Theorem 2.4. Hermite’s theory was made more precise by Julia (1917).
For more details about reduction theories of integral binary forms and polynomials of low degree we refer to Dickson, Vol. 3 (1919, reprinted 1971), Cremona (1999), Evertse and Győry (2017), Bhargava and Yang (2022), and for more general results and applications, also to Section 4 of the present paper and the references given there.
For the number of $\mathbb{Z}$-equivalence classes of cubic monic integral polynomials with given non-zero discriminant, no finiteness results were known before 1930. Then Delone and Nagell proved independently the following.
Theorem 2.5 (Delone, 1930; Nagell, 1930).
Up to $\mathbb{Z}$-equivalence, there are only finitely many irreducible cubic monic polynomials in $\mathbb{Z}[X]$ of given non-zero discriminant.
The proofs of Delone and Nagell of Theorem 2.5 were both ineffective, in that they did not provide a method to determine the polynomials. In fact, these proofs were based on a classical ineffective finiteness theorem of Thue (1909) on Thue equations, i.e. on equations of the form $F(x,y)=m$, $x,y\in\mathbb{Z}$, where $F\in\mathbb{Z}[X,Y]$ is a binary form of degree $\geq 3$ with discriminant $\neq 0$ and $m$ is a non-zero integer. In some concrete cases Delone and Faddeev (1940) made effective Theorem 2.5, and posed the problem to make it effective for any irreducible cubic monic polynomial. An effective version of Theorem 2.5 follows from the famous effective result of Baker (1968) on Thue equations.
3. Hermite’s attempt (1857) to extend the previous reduction results to polynomials of arbitrary degree
3.1. $\mathbf{GL_{n}(\mathbb{Z})}$-equivalence of decomposable forms
a
Hermite tried to extend his theorem (1851) on cubic integral binary forms resp. polynomials to the case of arbitrary degree $n\geq 4$, but without success. Instead, he proved a finiteness theorem with a weaker equivalence, see Theorem 3.2 below. Hermite’s notion of equivalence (called by us ‘Hermite equivalence’) is based on an equivalence relation for certain decomposable forms.
Consider decomposable forms of degree $n\geq 2$ in the same number $n$ of variables
$F(\mathbf{X})=a_{0}\prod_{i=1}^{n}(\alpha_{i,1}X_{1}+\cdots+\alpha_{i,n}X_{n})\in\mathbb{Z}[X_{1},\ldots,X_{n}],$
where $a_{0}$ is a non-zero rational number and $\alpha_{i,j}$ are algebraic numbers, not all zero, for $i,j=1,\ldots,n$. The discriminant of $F$ is defined as
$D(F):=a_{0}^{2}(\det(\alpha_{i,j}))^{2}.$
It is important to note that $D(F)$ is a rational integer.
Let $GL_{n}(\mathbb{Z})$ denote the multiplicative group of $n\times n$ integer matrices of determinant ±$\pm 1$. Two decomposable forms $F,G$ as above are called $GL_{n}(\mathbb{Z})$-equivalent if
$G(\mathbf{X})=\pm F(U\mathbf{X})\text{{ for some }}U\in GL_{n}(\mathbb{Z}),$
where $\mathbf{X}$ denotes the column vector of variables $(X_{1},\ldots,X_{n})^{T}$.
It is easy to see that two $GL_{n}(\mathbb{Z})$-equivalent decomposable forms in $n$ variables have the same discriminant.
Hermite proved the following.
Theorem 3.1 (Hermite, 1851).
Let $n$ and $D$ be integers with $n\geq 2$, $D\neq 0$. Then the decomposable forms in $\mathbb{Z}[X_{1},\ldots,X_{n}]$ of degree $n$ and discriminant $D$ lie in finitely many $GL_{n}(\mathbb{Z})$-equivalence classes.
3.2. Hermite equivalence of polynomials and Hermite’s finiteness theorem
a
Let
$f(X)=a_{0}(X-\alpha_{1})\cdots(X-\alpha_{n})\in\mathbb{Z}[X]$
be an integral polynomial with $a_{0}\in\mathbb{Z}\setminus\{0\}$, and $\alpha_{1},\ldots,\alpha_{n}\in\overline{\mathbb{Q}}$. Then the discriminant of $f$ is
$D(f)=a_{0}^{2n-2}\prod_{1\leq i<j\leq n}(\alpha_{i}-\alpha_{j})^{2}\in\mathbb{Z}.$
To $f$ we associate the decomposable form
$[f](\mathbf{X}):=a_{0}^{n-1}\prod_{i=1}^{n}(X_{1}+\alpha_{i}X_{2}+\cdots+\alpha_{i}^{n-1}X_{n})\in\mathbb{Z}[X_{1},\ldots,X_{n}].$
Using the properties of Vandermonde determinants, one can prove that
(3.1) $D([f])=D(f).$
The following equivalence relation was introduced by Hermite (1857):
  • Two polynomials $f,g\in\mathbb{Z}[X]$ of degree $n$ are said to be Hermite equivalent if the associated decomposable forms $[f]$ and $[g]$ are $GL_{n}(\mathbb{Z})$-equivalent, i.e.,
    $[g](\mathbf{X})=\pm[f](U\mathbf{X})\text{ {for some} }U\in GL_{n}(\mathbb{Z}).$
From (3.1) it follows directly that Hermite equivalent polynomials in $\mathbb{Z}[X]$ have the same discriminant.
Hermite’s Theorem 3.1 on decomposable forms and identity (3.1) imply the following finiteness theorem on polynomials.
Theorem 3.2 (Hermite, 1857).
Let $n\geq 2$ and $D\neq 0$ be integers. Then the polynomials $f\in\mathbb{Z}[X]$ of degree $n$ and of discriminant $D$ lie in finitely many Hermite equivalence classes.
Hermite’s proof is ineffective.
3.3. Comparison between Hermite equivalence and $GL_{2}(\mathbb{Z})$-equivalence and $\mathbb{Z}$-equivalence
a
In our five authors paper with Bhargava, Remete and Swaminathan (BEGyRS, 2023) we have integrated Hermite’s long-forgotten notion of equivalence and his finiteness theorem in the reduction theory, have corrected a faulty reference to Hermite’s result in Narkiewicz’ excellent book (2018) and compared Hermite’s theorem with the most significant results of this area; see the next Section 4.
In BEGyRS (2023) we proved that $GL_{2}(\mathbb{Z})$-equivalence and, in the monic case, $\mathbb{Z}$-equivalence imply Hermite equivalence.
Proposition 3.3 (BEGyRS, 2023).
Let $f,g\in\mathbb{Z}[X]$ be two $\mathbb{Z}$-equivalent, resp. $GL_{2}(\mathbb{Z})$-equivalent integral polynomials. Then they are Hermite equivalent.
Since $\mathbb{Z}$-equivalence implies $GL_{2}(\mathbb{Z})$-equivalence, it suffices to prove Proposition 3.3 for $GL_{2}(\mathbb{Z})$-equivalence. We recall the proof from BEGyRS (2023).
Proof.
Let $f,g$ in $\mathbb{Z}[X]$ be any two $GL_{2}(\mathbb{Z})$-equivalent polynomials. Then they can be written in the form $f(X)=a_{0}\prod_{i=1}^{n}(X-\alpha_{i})$ and $g(X)=\pm(cX+d)^{n}f\left(\frac{aX+b}{cX+d}\right)$, where $A:=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in GL_{2}(\mathbb{Z})$. Thus, we have
$g(X)=\pm a_{0}\prod_{i=1}^{n}(\beta_{i}X-\gamma_{i})\text{, where $\beta_{i}=c-a\alpha_{i},\,\gamma_{i}=-d+b\alpha_{i}$}$
for $i=1,\ldots,n$. Define the inner product of two column vectors
$\mathbf{x}=(x_{1},\ldots,x_{n})^{T},\,\,\mathbf{y}=(y_{1},\ldots,y_{n})^{T}\text{ by }\langle\mathbf{x},\mathbf{y}\rangle:=x_{1}y_{1}+\cdots+x_{n}y_{n}.$
Let as before $\mathbf{X}=(X_{1},\ldots,X_{n})^{T}$. Thus,
$[f](\mathbf{X})=a_{0}^{n-1}\prod_{i=1}^{n}\langle\mathbf{a}_{i},\mathbf{X}\rangle\text{, where }\mathbf{a}_{i}=(1,\alpha_{i},\ldots,\alpha_{i}^{n-1})^{T},$
$[g](\mathbf{X})$ =±$=\pm a_{0}^{n-1}\prod_{i=1}^{n}\langle\mathbf{b}_{i},\mathbf{X}\rangle\text{, where }\mathbf{b}_{i}=(\beta_{i}^{n-1},\beta_{i}^{n-2}\gamma_{i},\ldots,\gamma_{i}^{n-1})^{T}.$
Then $\mathbf{b}_{i}=t(A)\mathbf{a}_{i}$ with some $t(A)\in GL_{n}(\mathbb{Z})$ for $i=1,\ldots,n$. So
$[g](\mathbf{X})=$$\pm c^{n-1}\prod_{i=1}^{n}\langle t(A)\mathbf{a}_{i},\mathbf{X}\rangle=$
=±$=\pm c^{n-1}\prod_{i=1}^{n}\langle\mathbf{a}_{i},t(A)^{T}\mathbf{X}\rangle=\pm[f](t(A)^{T}\mathbf{X}),$
i.e. $f$ and $g$ are indeed Hermite equivalent. ∎
For integral polynomials of degree $2$ and $3$, Hermite equivalence and $GL_{2}(\mathbb{Z})$-equivalence coincide. For quadratic polynomials this is trivial, while for cubic polynomials this follows from a result of Delone and Faddeev (1940).
In BEGyRS (2023) we gave, for every $n\geq 4$ and both for the nonmonic and for the monic case, infinite collections of polynomials in $\mathbb{Z}[X]$ with degree $n$ that are Hermite equivalent but not $GL_{2}(\mathbb{Z})$-equivalent. More precisely we proved the following.
Theorem 3.4 (BEGyRS, 2023).
Let $n$ be an integer $\geq 4$.
  1. (i)
    There exist infinitely many Hermite equivalence classes of properly nonmonic [^1] That is, not $GL_{2}(\mathbb{Z})$-equivalent to any monic polynomial irreducible polynomials of degree $n$ that split into more than one $GL_{2}(\mathbb{Z})$-equivalence class.
  2. (ii)
    There exist infinitely many Hermite equivalence classes of monic irreducible polynomials of degree $n$ that split into more than one $GL_{2}(\mathbb{Z})$-equivalence class.
In the monic case every $GL_{2}(\mathbb{Z})$-class contains a $\mathbb{Z}$-equivalence class, hence in (ii) $GL_{2}(\mathbb{Z})$-equivalence can be replaced by $\mathbb{Z}$-equivalence.
We proved Theorem 3.4 simultaneously for the cases (i) and (ii). We constructed, for every integer $n\geq 4$, an infinite parametric family of pairs $(f_{t,c}^{(n)},g_{t,c}^{(n)})$ of primitive  [^2] An integral polynomial is called primitive if its coefficients have greatest common divisor $1$, irreducible polynomials $f_{t,c}^{(n)}$, $g_{t,c}^{(n)}$ of degree $n$, where $c$ runs through $1$ and an infinite set of primes, and $t$ runs through an infinite set of primes with $t\neq c$ with the following properties:
(3.2) $f_{t,c}^{(n)},\,g_{t,c}^{(n)}$ have leading coefficient $c$ and are properly
nonmonic if $c>1$;
(3.3) $f_{t,c}^{(n)},\,g_{t,c}^{(n)}$ are Hermite equivalent;
(3.4) $f_{t,c}^{(n)},\,g_{t,c}^{(n)}$ are not $GL_{2}(\mathbb{Z})$-equivalent;
(3.5) the pairs $(f_{t,c}^{(n)},g_{t,c}^{(n)})$ lie in different Hermite equivalence
classes.
The main steps of the proof are as follows. From the construction of $f_{t,c}^{(n)}$ and $g_{t,c}^{(n)}$ it is easy to show that (3.2) and (3.3) hold. The proof of (3.4) is more complicated. It requires the use of an irreducibility theorem of Dumas (1906), Chebotarev’s density theorem, and Dirichlet’s theorem on primes in arithmetic progressions. Finally, $f_{t,c}^{(n)}$ is so chosen that if we fix $n,c$ and let $t\to\infty$ then the absolute value of the discriminant of $f_{t,c}^{(n)}$ tends to $\infty$. Since Hermite equivalent polynomials have the same discriminant, the pairs $(f_{t,c}^{(n)},g_{t,c}^{(n)})$ lie in infinitely many different Hermite equivalence classes, that is, (3.5) follows.
Remark.
We note that in our paper BEGyRS (2023) it turned out that the Hermite equivalence class of a polynomial has a very natural interpretation in terms of the so-called invariant ring and invariant ideal associated with the polynomial, see Appendix B for more details. This fact turned out to be important in the above proofs.
Proposition 3.3 and Theorem 3.4 imply that $GL_{2}(\mathbb{Z})$-equivalence, resp. $\mathbb{Z}$-equivalence are stronger than Hermite equivalence, and hence that Hermite’s Theorem 3.2 is much weaker than the most significant results of this area presented in Section 4 below.
4. Reduction theory of integral polynomials of given non-zero discriminant: the general case
As we mentioned in the Introduction, the breakthroughs in the reduction theory due to Birch and Merriman (1972), Győry (1973), and Evertse and Győry (1991) settled the old problem of Hermite (1857), to prove that for every given $n\geq 2$ and $D\not=0$ there are up to $GL_{2}(\mathbb{Z})$-equivalence only finitely many polynomials $f\in\mathbb{Z}[X]$ of degree $n$ and discriminant $D$, and to determine these effectively. We state the results in more detail.
4.1. The theorems of Birch and Merriman (1972), Győry (1973) and Evertse and Győry (1991)
a
Theorem 4.1 (Birch and Merriman, 1972).
Let $n\geq 2$ and $D\neq 0$. There are only finitely many $GL_{2}(\mathbb{Z})$-equivalence classes of polynomials in $\mathbb{Z}[X]$ of degree $n$ and discriminant $D$.
Birch and Merriman established this theorem in an equivalent form, in terms of integral binary forms. Their proof uses the finiteness of the number of solutions of unit equations $ax+by=1$ in units $x,y$ of the ring of integers of a number field, for which at the time effective proofs were available, but it combines this with some ineffective arguments. Consequently, Birch’s and Merriman’s proof of Theorem 4.1 is ineffective.
For monic polynomials, the corresponding result with $\mathbb{Z}$-equivalence was proved independently by Győry (1973) but in an effective form. This proved to be of crucial importance in many applications; see e.g. Sections 5 to 8 below and Evertse and Győry (2017).
Theorem 4.2 (Győry, 1973).
Let $f\in\mathbb{Z}[X]$ be a monic polynomial of degree $n\geq 2$ with discriminant $D\neq 0$. Then
  1. (i)
    $n\leq c_{1}(|D|)$, and
  2. (ii)
    there is a monic $g\in\mathbb{Z}[X]$, $\mathbb{Z}$-equivalent to $f$, such that
    $H(g)\leq c_{2}(n,|D|),$
where $c_{1}$ and $c_{2}$ are effectively computable positive numbers depending on $D$, resp. on $n$ and $|D|$.
Corollary 4.3 (Győry, 1973).
There are only finitely many $\mathbb{Z}$-equivalence classes of monic polynomials in $\mathbb{Z}[X]$ of given non-zero discriminant, and a full set of representatives of these classes can be effectively determined.
In Győry (1974), an explicit version was given; see below.
In his proof of Theorem 4.2, Győry combined his own effective result on unit equations obtained by Baker’s method, with his so-called ‘graph method’. We sketch below the proof of Theorem 4.2.
Theorem 4.1, resp. Theorem 4.2 and its Corollary 4.3 generalized the corresponding results presented in Section 2 to polynomials of any degree $n\geq 3$; Theorem 4.1 gives an ineffective generalization and Theorem 4.2 an effective generalization.
In 1991, Evertse and Győry gave a new, effective proof for Birch’s and Merriman’s theorem, proving the following.
Theorem 4.4 (Evertse and Győry, 1991a).
Let $f\in\mathbb{Z}[X]$ be a polynomial of degree $n\geq 2$ and discriminant $D\neq 0$. There is a $g\in\mathbb{Z}[X]$, $GL_{2}(\mathbb{Z})$-equivalent to $f$, such that
$H(g)\leq c_{3}(n,|D|),$
where $c_{3}(n,|D|)$ is an effectively computable number, given explicitly in terms of $n$ and $|D|$.
As was mentioned above, Theorems 4.2 and 4.4 provided an effective version of the general reduction theory of integral polynomials of given non-zero discriminant.
The main tool in our proof of Theorem 4.4 is an effective result of Győry (1979) on homogeneous unit equations in three unknowns, whose proof is also based on Baker’s theory of logarithmic forms.
Theorems 4.2 and 4.4, their explicit versions below and their various generalizations have a great number of consequences and applications; see our book Evertse and Győry (2017) and Sections 5 to 8 below.
4.2. Explicit versions of theorems of Győry (1973) and Evertse and Győry (1991a)
a
First we present explicit versions of Theorem 2.1, Theorem 2.2 and Theorem 2.4 in the quadratic and cubic cases. An explicit version of Theorem 2.1 is the following.
Theorem 2.1*.
Let $f\in\mathbb{Z}[X]$ be a quadratic polynomial of discriminant $D\neq 0$. Then $f$ is $GL_{2}(\mathbb{Z})$-equivalent to a quadratic polynomial $g\in\mathbb{Z}[X]$ such that
  1. (i)
    $H(g)\leq|D|/3$ if $D<0$;
  2. (ii)
    $H(g)\leq|D|/4$ if $D>0$ and $f$ is irreducible;
  3. (iii)
    $H(g)\leq D^{1/2}$ if $D>0$ and $f$ is reducible.
In the cubic case, we have the following.
Theorem 2.4*.
Let $f\in\mathbb{Z}[X]$ be a cubic polynomial of discriminant $D\neq 0$. Then $f$ is $GL_{2}(\mathbb{Z})$-equivalent to a cubic polynomial $g\in\mathbb{Z}[X]$ such that
  1. (i)
    $H(g)\leq\frac{64}{27}|D|^{1/2}$ if $f$ is irreducible;
  2. (ii)
    $H(g)\leq\frac{64}{3\sqrt{3}}|D|$ if $f$ is reducible.
We note that the arguments in the proofs of Theorems 2.1* and 2.4* are a variation on the arguments in Julia (1917). For the details we refer to Subsection 13.1 of the book of Evertse and Győry (2017).
In the monic case, it is relatively simple to prove the following explicit version of Theorem 2.2.
Theorem 2.2*.
For any monic quadratic polynomial $f\in\mathbb{Z}[X]$ with discriminant $D\neq 0$, there exist $g\in\mathbb{Z}[X]$, $\mathbb{Z}$-equivalent to $f$, such that
$H(g)\leq|D|/4+1.$
As was mentioned above, the first explicit version of Theorem 4.2 was given in Győry (1974). This was improved in 2017 by the authors.
Theorem 4.2* (Evertse and Győry, 2017).
Let $f\in\mathbb{Z}[X]$ be a monic polynomial of degree $n\geq 2$ and discriminant $D\neq 0$. Then $f$ is $\mathbb{Z}$-equivalent to a polynomial $g\in\mathbb{Z}[X]$ for which
(4.1) $H(g)\leq\exp\{n^{20}8^{n^{2}+19}(|D|(\log|D|)^{n})^{n-1}\}.$
A completely explicit, improved version of Theorem 4.4 was also established by the authors.
Theorem 4.4* (Evertse and Győry, 2017).
Let $f\in\mathbb{Z}[X]$ be a polynomial of degree $n\geq 2$ and discriminant $D\neq 0$. Then $f$ is $GL_{2}(\mathbb{Z})$-equivalent to a polynomial $g\in\mathbb{Z}[X]$ for which
(4.2) $H(g)\leq\exp\{(4^{2}n^{3})^{25n^{2}}\cdot|D|^{5n-3}\}.$
In both Theorems 4.2* and 4.4*, the degree $n$ of $f$ can also be explicitly estimated from above in terms of $|D|$.
Theorem 4.5 (Győry, 1974).
Every polynomial $f\in\mathbb{Z}[X]$ with discriminant $D\neq 0$ has degree at most
$3+2\log|D|/\log 3$
For monic polynomials $f\in\mathbb{Z}[X]$, the upper bound can be improved slightly to $2+2\log|D|/\log 3$.
Theorem 4.4 together with Theorem 4.5 implies the following analogue of Corollary 4.3.
Corollary 4.6 (Evertse and Győry, 1991).
There are only finitely many $GL_{2}(\mathbb{Z})$-equivalence classes of polynomials in $\mathbb{Z}[X]$ of given non-zero discriminant, and a full set of representatives of these classes can be effectively determined.
As was mentioned above, for $n\geq 4$ resp. $n\geq 3$ the proofs of Theorems 4.2, 4.4, 4.2* and 4.4* are based on effective results of Győry on unit equations whose proofs depend on Baker’s theory of logarithmic forms. The exponential feature of the bounds in (4.1) and (4.2) is a consequence of the use of Baker’s method. It is likely that the bounds in (4.1) and (4.2) can be replaced by some bounds polynomial in terms of $|D|$. This can be achieved if we restrict ourselves to polynomials $f\in\mathbb{Z}[X]$ having a fixed splitting field $G$ over $\mathbb{Q}$. In this case the bounds in (4.1) and (4.2) can be replaced by bounds of the form
$c_{4}(n,G)|D|^{c_{5}(n,G)},$
where $c_{4}(n,G),c_{5}(n,G)$ are effectively computable numbers which depend only on $n$ and the discriminant of $G$; see Győry (1984, 1998) resp. Evertse and Győry (1991a). The following conjecture seems plausible.
Conjecture 4.7.
Let $f\in\mathbb{Z}[X]$ be a polynomial resp. a monic polynomial of degree $n\geq 4$ resp. $n\geq 3$ with discriminand $D\neq 0$. Then $f$ is $GL_{2}(\mathbb{Z})$-equivalent resp. $\mathbb{Z}$-equivalent to a polynomial resp. monic polynomial such that
$H(g)\leq c_{6}(n)|D|^{c_{7}(n)}$
where $c_{6}(n),c_{7}(n)$ depend only on $n$.
The first part of the conjecture is formulated in Section 15 of Evertse and Győry (2017).
Evertse proved the following what one may call semi-effective result.
Theorem 4.8 (Evertse, 1993).
Let $f\in\mathbb{Z}[X]$ be a polynomial of degree $n\geq 4$ and of discriminant $D\neq 0$, having splitting field $G$ over $\mathbb{Q}$. Then $f$ is $GL_{2}(\mathbb{Z})$-equivalent to a polynomial $g$ of height
$H(g)\leq c_{8}(n,G)|D|^{21/n}.$
Here $c_{8}(n,G)$ is a number depending only on $n$ and $G$, which is not effectively computable by the method of proof. Evertse’ s proof is based on a version of Roth’s Diophantine approximation theorem over number fields, which is ineffective.
4.3. The general approach of Győry (1973) for proving Theorem 4.2
a
For the complete proof we refer to Győry (1973) or the Subsections 6.6 and 6.7 of Evertse and Győry (2017). Here, we give a brief outline. The primitive minimal polynomial $f_{\alpha}$ of an algebraic number $\alpha$ is the minimal polynomial of $\alpha$ whose coefficients are integers with greatest common divisor $1$ and whose leading coefficient is positive. The height $H(\alpha)$ of an algebraic integer $\alpha$ is the height of $f_{\alpha}$.
The proof of the boundedness of the degree $n$ of $f$ in terms of $D$ uses Minkowski’s lower bound for the absolute value of the discriminant of a number field in terms of its degree and combines this with an elementary argument. Henceforth, we restrict ourselves to the case that $f\in\mathbb{Z}[X]$ is a monic irreducible polynomial of discriminant $D\neq 0$ and of fixed degree $n$ with $2\leq n\leq c_{1}(|D|)$.
Thus, let $f\in\mathbb{Z}[X]$ be such a polynomial. Then the main steps of the proof of Theorem 4.2 are as follows.
  1. (1)
    Denote by $\alpha_{1},\ldots,\alpha_{n}$ the zeros of $f$, and by $G$ the splitting field of $f$ over $\mathbb{Q}$. Then $[G:\mathbb{Q}]\leq n!$ and the absolute value $|D_{G}|$ of the discriminant of $G$ can be estimated from above by a constant $c_{9}(D)$. Here and below $c_{9},\ldots$ are effectively computable numbers depending only on $D$.
  2. (2)
    Putting $\Delta_{ij}:=\alpha_{i}-\alpha_{j}$ we have
    $\prod_{1\leq i<j\leq n}\Delta_{ij}^{2}=D,$
    which implies $|N_{G/\mathbb{Q}}\Delta_{ij}|\leq c_{10}(D)$. It follows that
    (4.3) $\Delta_{ij}=\delta_{ij}\varepsilon_{ij}\text{, where }H(\delta_{ij})\leq c_{11}(D)$
    and $\varepsilon_{ij}$ is a unit in the ring of integers of $G$.
  3. (3)
    The following identity plays a basic role in the proof:
    (4.4) $\Delta_{ij}+\Delta_{jk}=\Delta_{ik}\text{ for every }i,j,k.$
    Consider the graph, whose vertices are $\Delta_{ij}$ $(1\leq i\not=j\leq n)$ and whose edges are $[\Delta_{ij},\Delta_{ik}]$, $[\Delta_{ij},\Delta_{jk}]$ ($1\leq i\not=j\leq n$, $k\not=i,j$). This graph is obviously connected.
  4. (4)
    Equations (4.3) and (4.4) give rise to a ‘connected’ system of unit equations
    (4.5) $\delta_{ijk}\varepsilon_{ijk}+\tau_{ijk}\nu_{ijk}=1,$
    where $\delta_{ijk}:=\delta_{ij}/\delta_{ik}$, $\tau_{ijk}:=\delta_{jk}/\delta_{ik}$ are non-zero elements of $G$ with heights effectively bounded above in terms of $|D|$ only, and $\varepsilon_{ijk}:=\varepsilon_{ij}/\varepsilon_{ik}$, $\nu_{ijk}:=\varepsilon_{jk}/\varepsilon_{ik}$ are unknown units in the ring of integers of $G$.
  5. (5)
    Represent $\varepsilon_{ijk}$ as
    $\varepsilon_{ijk}=\zeta_{ijk}\varrho_{1}^{a_{ijk,1}}\cdots\varrho_{r}^{a_{ijk},r},$
    and similarly $\nu_{ijk}$, where $\zeta_{ijk}$ is a root of unity, $\varrho_{1},\ldots,\varrho_{r}$ a fundamental system of units in $G$ with heights effectively bounded above in terms of $|D|$ and, by Dirichlet theorem $r\leq n!-1$.
  6. (6)
    Applying Baker’s method to (4.5), we get effective upper bounds for $|a_{ijk,1}|,\ldots,|a_{ijk},r|$ that depend only on $|D|$. In view of (4.3), this leads to upper bounds for the heights of the quotients $\Delta_{ij}/\Delta_{ik}=\delta_{ijk}\varepsilon_{ijk}$ for each triple $\{i,j,k\}\subset\{1,\ldots,n\}$, depending on $G,n$ and $|D|$, and so eventually only on $|D|$, and likewise for $\Delta_{jk}/\Delta_{ik}$.
    Remark: in Győry (1974), this was the first application of Baker’s method to general unit equations of the form (4.5) with explicit bounds.
  7. (7)
    Using the connectedness of the unit equations involved, this yields effective upper bounds for the height of $\Delta_{ij}$ for every $i,j$, depending only on $|D|$. Indeed, one first obtains an upper bound for the height of any quotient $\Delta_{ij}/\Delta_{kl}$ via
    $\frac{\Delta_{ij}}{\Delta_{kl}}=\frac{\Delta_{ij}}{\Delta_{ik}}\cdot\frac{\Delta_{ik}}{\Delta_{kl}}$
    (using the path $\Delta_{ij}\to\Delta_{ik}\to\Delta_{kl}$ in the graph) and subsequently for the height of each $\Delta_{ij}$ via
    $\Delta_{ij}^{n(n-1)}=\pm D\cdot\prod_{1\leq k\not=l\leq n}\frac{\Delta_{ij}}{\Delta_{kl}}.$
  8. (8)
    Adding the differences $\Delta_{ij}=\alpha_{i}-\alpha_{j}$ for fixed $i$ and for $j=1,\ldots,n$, using the fact that $\alpha_{1}+\cdots+\alpha_{n}\in\mathbb{Z}$, putting $\alpha_{1}+\cdots+\alpha_{n}=na+a^{\prime}$ with $a,a^{\prime}\in\mathbb{Z}$, $0\leq a^{\prime}<n$, and writing
    $\beta_{i}:=\alpha_{i}-a\text{ for }i=1,\ldots,n,$
    $g(X)=\prod_{i=1}^{n}(X-\beta_{i}),$
    we have that $g(X)=f(X+a)\in\mathbb{Z}[X]$ and that the height of $g$ has an effective upper bound depending only on $|D|$. ∎
We note that for cubic and quartic monic polynomials $f\in\mathbb{Z}[X]$ of given non-zero discriminant Klaska (2021, 2022) devises another approach for proving Corollary 4.3 via the theory of integral points on elliptic curves.
4.4. A brief sketch of the proof of a less precise version of Theorem 4.4
a
Take an integral polynomial $f\in\mathbb{Z}[X]$ of degree $n$ and discriminant $D\neq 0$. In view of Theorems 2.1, 2.2 and 2.4 we may assume that $n\geq 4$. The discriminant of the splitting field of $f$ can be estimated from above in terms of $D$, and by the Hermite–Minkowski Theorem, this leaves only a finite, effectively determinable collection of possible splitting fields for $f$. So we may restrict ourselves to polynomials $f$ with given splitting field $G$ and ring of integers $\mathcal{O}_{G}$.
Take such $f$ and pick a factorization of $f$,
(4.6) $f=\prod_{i=1}^{n}(\alpha_{i}X-\beta_{i})\,\,\text{over }\overline{\mathbb{Q}},$
such that the number of linear factors with real coefficients is maximal, and the factors with complex coefficients fall apart into complex conjugate pairs. After multiplying $f$ by a small positive rational integer, which can be effectively bounded in terms of $G$, hence in terms of $D$ and which is negligible compared with the other estimates arising from the application of Baker’s method, we may assume that $f$ has such a factorization with $\alpha_{i},\beta_{i}\in\mathcal{O}_{G}$ for $i=1,\ldots,n$. Put
$\Delta_{ij}:=\alpha_{i}\beta_{j}-\alpha_{j}\beta_{i}\ \ \text{for $1\leq i,j\leq n$.}$
We now follow the approach of Evertse and Győry (2017), chapters 13 and 14. We outline the main steps of the proof.
  1. (1)
    We start with a small variation on the reduction theory of Hermite (1848, 1851) and Julia (1917). Let ${\bf t}=(t_{1},\ldots,t_{n})$ be a tuple of positive reals such that $t_{i}=t_{j}$ for each pair $(i,j)$ such that $\alpha_{i},\beta_{i}$ are the complex conjugates of $\alpha_{j},\beta_{j}$. Consider the positive definite quadratic form
    $\Phi_{f,{\bf t}}(X,Y):=\sum_{i=1}^{n}t_{i}^{-2}(\alpha_{i}X-\beta_{i}Y)(\overline{\alpha_{i}}X-\overline{\beta_{i}}Y).$
    By Gauss’ reduction theory for positive definite binary quadratic forms, there is $\big{(}\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big{)}\in GL_{2}(\mathbb{Z})$ such that $\Phi_{f,{\bf t}}(aX+bY,cX+dY)$ is reduced, i.e., equal to $AX^{2}+BXY+CY^{2}$ with $|B|\leq A\leq C$. Define the polynomial
    $g(X)=(cX+d)^{n}f\left(\frac{aX+b}{cX+d}\right),$
    which is $GL_{2}(\mathbb{Z})$-equivalent to $f$. We denote by $H(g)$ the height of $g$. We recall Theorem 13.1.3 of Evertse and Győry (2017), and refer for the elementary proof to section 13.1 of that book.
    Proposition 4.9.
    Let
    $M:=t_{1}\cdots t_{n},\ \ R:=\left(\sum_{1\leq i<j\leq n}\frac{|\Delta_{ij}|^{2}}{t_{i}^{2}t_{j}^{2}}\right)^{2}.$
    Then
    $H(g)\leq\left(\frac{4}{n\sqrt{3}}\right)^{n}M^{2}R^{n}$
    if $f$ has no root in $\mathbb{Q}$, and
    $H(g)\leq\left(\frac{2}{\sqrt{n}}\right)^{n}\cdot\left(\frac{2}{\sqrt{3(n-1)}}\right)^{n(n-1)/(n-2)}(M^{2}R^{n})^{(n-1)/(n-2)}$
    if $f$ does have a root in $\mathbb{Q}$.
  2. (2)
    For any quadruple $i,j,k,l$ of distinct indices we have the identity
    (4.7) $\Delta_{ij}\Delta_{kl}+\Delta_{jk}\Delta_{il}=\Delta_{ik}\Delta_{jl}.$
    Notice that all terms $\Delta_{ij}$ are in $\mathcal{O}_{G}$ and divide $D$. Hence $|N_{G/\mathbb{Q}}(\Delta_{ij})|\leq|D|^{[G:\mathbb{Q}]}$ for all $i,j$ where $[G:\mathbb{Q}]\leq n!$. As above in Section 4.3, we can express each term $\Delta_{ij}$ as a product of an element of height effectively bounded in terms of $n,D$ and a unit from $\mathcal{O}_{G}$. By substituting this into the identities (4.7) we obtain homogeneous unit equations in three terms. Dividing (4.7) by $\Delta_{ik}\Delta_{jl}$ we get unit equations like in (4.5) above, and using the theorem of Győry (1974) we obtain effective upper bounds for the heights of the quotients $\Delta_{ij}\Delta_{kl}/\Delta_{ik}\Delta_{jl}$.
  3. (3)
    To obtain an effective upper bound for the height of $g$ in terms of $D$, it suffices to effectively estimate the quantities $M$ and $R$ from Proposition 4.9 from above in terms of $D$, for a suitable choice of the $t_{i}$. For the $t_{i}$ we choose
    $t_{i}:=\left(\prod_{k=1,k\not=i}^{n}|\Delta_{ik}|\right)^{1/(n-2)}\ \ \text{for }i=1,\ldots,n.$
    With this choice,
    $M=|D|^{1/(n-2)}$
    and
    $\frac{|\Delta_{ij}|}{t_{i}t_{j}}=\left(|D|^{-1}\cdot\prod_{k,l}\left|\frac{\Delta_{ij}\Delta_{kl}}{\Delta_{ik}\Delta_{jl}}\right|\right)^{1/(n-1)(n-2)},$
    where the product is taken over all pairs of indices $k,l$ such that $1\leq k,l\leq n$, $k\not=i,j$, $l\not=i,j$ and $k\not=l$. By inserting the upper bounds for the heights of the quantities $\Delta_{ij}\Delta_{kl}/\Delta_{ik}\Delta_{jl}$ obtained in the previous step, we can estimate from above $M$ and $R$, and subsequently $H(g)$, effectively in terms of $D$ only. ∎
5. Consequences of Theorem 4.2 in algebraic number theory, and in particular for monogenicity of number fields and their orders
Theorem 4.2 and its variants in Győry (1973, 74, 76, 78a,b) led to breakthroughs in the effective theory of number fields. They furnished general effective finiteness results for integral elements of given non-zero discriminant resp. given index in any number field $K$ and its orders. In particular, in Győry (1976), Theorem 4.2 provided the first general algorithm for deciding the monogenicity of $K$ and, more generally, of any order of $K$, and for determining all power integral bases in $K$ and in its orders. For later generalizations, applications and comprehensive treatment of this extensive area, we refer to Győry (1980a, 1980b, 2000), Evertse and Győry (2017), the references given there, and to Section 8 of the present paper.
For convenience, we present the general effective finiteness theorems mentioned above in their simplest form. Further, we briefly sketch how to deduce the consequences directly from Theorem 4.2 via its Corollary 5.1.
5.1. Preliminaries
a
Throughout this section, $K$ will denote a number field of degree $n\geq 2$ with ring of integers $\mathcal{O}_{K}$ and discriminant $D_{K}$. For a primitive integral element $\alpha$ of $K$ we denote by $f_{\alpha}(X)$ the minimal (monic) polynomial of $\alpha$ in $\mathbb{Z}[X]$. Thus, $f_{\alpha}(X)=\prod_{i=1}^{n}(X-\alpha^{(i)})$, where $n$ is the degree of $\alpha$ and $\alpha^{(1)}=\alpha,\alpha^{(2)},\ldots,\alpha^{(n)}$ denote the conjugates of $\alpha$. Then the discriminant of $\alpha$ relative to the extension $K/\mathbb{Q}$ is defined by
(5.1) $D_{K/\mathbb{Q}}(\alpha):=D(f_{\alpha})=\prod_{1\leq i<j\leq n}(\alpha^{(i)}-\alpha^{(j)})^{2}.$
Let now $\mathcal{O}$ be an order of $K$ (i.e., a subring of $K$ that as a $\mathbb{Z}$-module is free of rank $n=[K:\mathbb{Q}]$), and $D_{\mathcal{O}}$ its discriminant. Then $\mathcal{O}$ is a subring of $\mathcal{O}_{K}$. For $\alpha\in\mathcal{O}_{K}$ resp. $\alpha\in\mathcal{O}$, we denote by $I(\alpha):=[\mathcal{O}_{K}:\mathbb{Z}[\alpha]]$ resp. by $I_{\mathcal{O}}(\alpha):=[\mathcal{O}:\mathbb{Z}[\alpha]]$ the index of $\alpha$ in $\mathcal{O}_{K}$ resp. in $\mathcal{O}$. Then, as is known,
(5.2a) $D_{K/\mathbb{Q}}(\alpha)=I^{2}(\alpha)D_{K}\text{ if }\alpha\in\mathcal{O}_{K},$
and
(5.2b) $D_{K/\mathbb{Q}}(\alpha)=I^{2}_{\mathcal{O}}(\alpha)D_{\mathcal{O}}\text{ if }\alpha\in\mathcal{O},$
where $D_{\mathcal{O}}$ is the discriminant of $\mathcal{O}$, i.e., of any $\mathbb{Z}$-module basis of $\mathcal{O}$.
Two algebraic integers $\alpha$, $\beta$ are called $\mathbb{Z}$-equivalent if $\beta=\pm\alpha+a$ for some $a\in\mathbb{Z}$. If $\alpha$ and $\beta$ are $\mathbb{Z}$-equivalent then so are $f_{\alpha}$ and $f_{\beta}$. Conversely, if $f_{\alpha}$ and $f_{\beta}$ are $\mathbb{Z}$-equivalent then $\alpha$ is $\mathbb{Z}$-equivalent to a conjugate of $\beta$.
Clearly, $\mathbb{Z}$-equivalence elements in $\mathcal{O}_{K}$ resp. $\mathcal{O}$ have the same discriminant and hence the same index in $\mathcal{O}_{K}$ resp. in $\mathcal{O}$.
A number field $K$ is called monogenic if $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ for some $\alpha\in\mathcal{O}_{K}$. This is equivalent to the fact that $I(\alpha)=1$ and that $\{1,\alpha,\ldots,\alpha^{n-1}\}$ is a power integral basis in $K$, i.e., a $\mathbb{Z}$-module basis of $\mathcal{O}_{K}$. Similarly, an order $\mathcal{O}$ of $K$ is said to be monogenic if $\mathcal{O}=\mathbb{Z}[\alpha]$, i.e. if $I_{\mathcal{O}}(\alpha)=1$ for some $\alpha\in\mathcal{O}$. Clearly, if $\mathcal{O}=\mathbb{Z}[\alpha]$ then also $\mathcal{O}=\mathbb{Z}[\beta]$ for every $\beta$ that is $\mathbb{Z}$-equivalent to $\alpha$.
Further, $K$ resp. $\mathcal{O}$ is called $k\,(\geq 1)$ times monogenic if $\mathcal{O}_{K}$ resp. $\mathcal{O}$ equals $\mathbb{Z}[\alpha_{1}]=\cdots=\mathbb{Z}[\alpha_{k}]$ for some pairwise $\mathbb{Z}$-inequivalent $\alpha_{1},\ldots,\alpha_{k}$ in $\mathcal{O}_{K}$ resp. in $\mathcal{O}$. In case that in the above definition $k$ is maximal, it is called the multiplicity of the monogenicity of $K$, resp. $\mathcal{O}$.
5.2. Most important consequences of Theorem 4.2 in number fields
a
By the height $H(\alpha)$ of an algebraic integer $\alpha$ we mean the height $H(f_{\alpha})$. Further, the discriminant $D(\alpha)$ of $\alpha$ is defined as $D(\alpha):=D(f_{\alpha})$.
Corollary 5.1 (of Theorem 4.2).
Let $\alpha$ be an algebraic integer of degree $n\geq 2$ and discriminant $D\neq 0$. Then
  1. (i)
    $n\leq c_{1}(|D|)$, and
  2. (ii)
    there is an algebraic integer $\beta$, $\mathbb{Z}$-equivalent to $\alpha$, such that
$H(\beta)\leq c_{2}(n,|D|),$
where $c_{1},c_{2}$ denote the same effectively computable positive numbers as in Theorem 4.2.
This implies that there are only finitely many $\mathbb{Z}$-equivalence classes of algebraic integers with a given non-zero discriminant, and that a full set of representatives of these classes can be effectively determined.
Corollary 5.1 was proved in Győry (1973) as ‘Corollaire 3’ of the ‘Théorème’, the main result of the paper. This finiteness result in an ineffective form follows also from the work of Birch and Merriman (1972), which was independent of Győry (1973). Corollary 5.1 confirmed in full generality and in effective form a conjecture of Nagell (1967). Previously, the cubic case was settled independently by Delone (1930) and Nagell (1930), and the quartic case by Nagell (1967).
Finally, we note that Corollary 5.1 easily follows from Theorem 4.2. Indeed, if $\alpha$ is an algebraic integer with the properties specified in Corollary 5.1, then by (5.1), $D(f_{\alpha})=D$ and $\deg f_{\alpha}=n$. Further, by Theorem 4.2 $f_{\alpha}$ is $\mathbb{Z}$-equivalent to some monic $g\in\mathbb{Z}[X]$ with degree $n$ and discriminant $D$ such that $n\leq c_{1}(|D|)$ and $H(g)\leq c_{2}(n,|D|)$, where $c_{1},c_{2}$ denote the effectively computable numbers occurring in Theorem 4.2. But then $\alpha$ is $\mathbb{Z}$-equivalent to a zero of $g$, say $\beta$, whence $\deg\beta\leq c_{1}(|D|)$ and $H(\beta)\leq c_{2}(n,|D|)$ follow.∎
We note that Corollary 5.1 above and also Corollaire 1 of Győry (1973) imply in an effective way that there are only finitely many algebraic units in $\overline{\mathbb{Q}}$ with given discriminant. This gave the effective solution to Problem 19 in the book Narkiewicz (1974).
Consider again the above number field $K$ of degree $n$ with ring of integers $\mathcal{O}_{K}$ and discriminant $D_{K}$, and let $\mathcal{O}$ be an order of $K$ with discriminant $D_{\mathcal{O}}$. Restricting ourselves in Corollary 5.1 to the integers in $\mathcal{O}$, we get the following.
Corollary 5.2 (of Theorem 4.2).
Let $\mathcal{O}$ be an order of $K$, and let $\alpha\in\mathcal{O}$ be a primitive element of $K$ with index $I_{\mathcal{O}}$ in $\mathcal{O}$. Then $\alpha$ is $\mathbb{Z}$-equivalent to some $\beta$ in $\mathcal{O}$ such that
$H(\beta)\leq c_{2}(n,I^{2}_{\mathcal{O}}\cdot|D_{\mathcal{O}}|).$
If in particular $\mathcal{O}=\mathcal{O}_{K}$ and $\alpha\in\mathcal{O}_{K}$ with index $I$ in $\mathcal{O}_{K}$, then the same holds with $I_{\mathcal{O}},\,D_{\mathcal{O}}$ replaced by $I$ and $D_{K}$ respectively, where $c_{2}$ denotes the effectively computable number occurring in Theorem 4.2.
This implies that there are only finitely many $\mathbb{Z}$-equivalence classes of $\alpha$ in $\mathcal{O}$ resp. in $\mathcal{O}_{K}$ with given index $I_{\mathcal{O}}$ in $\mathcal{O}$ resp. $I$ in $\mathcal{O}_{K}$, and a full set of representatives of these classes can be effectively determined.
Corollary 5.2 follows immediately from Corollary 5.1, using (5.2b) resp. (5.2a). It was proved in Győry (1976) in an effective and quantitative form.
Corollary 5.3 below is the most influential consequence of Theorem 4.2. It gave the first general algorithm for deciding the monogenicity, and multiplicity of monogenicity of $K$ and its orders $\mathcal{O}$, and for determining all power integral bases in $K$ and its orders.
Corollary 5.3 (of Theorem 4.2).
If $\alpha$ is an element of $\mathcal{O}$ resp. of $\mathcal{O}_{K}$ such that $\mathcal{O}=\mathbb{Z}[\alpha]$ resp. $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ then $\alpha$ is $\mathbb{Z}$-equivalent to some $\beta$ in $\mathcal{O}$ resp. in $\mathcal{O}_{K}$ such that its height $H(\beta)$ is at most $c_{2}(n,|D_{\mathcal{O}}|)$ resp. $c_{2}(n,|D_{K}|)$, where $c_{2}$ denotes the effectively computable expression in Theorem 4.2 with $|D|$ replaced by $|D_{\mathcal{O}}|$ resp. by $|D_{K}|$.
It follows from Corollary 5.3 that there are only finitely many $\mathbb{Z}$-equivalence classes of $\alpha$ in $\mathcal{O}$ resp. in $\mathcal{O}_{K}$ such that $\mathcal{O}=\mathbb{Z}[\alpha]$ resp. $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$, and a full set of representatives of these classes can be effectively determined.
Corollary 5.3 is an immediate consequence of Corollary 5.2, choosing $I_{\mathcal{O}}=1$ resp. $I=1$. It was established in Győry (1976) with explicit expressions for $c_{1}$ and $c_{2}$.
Remark.
With the above formulation of Corollaries 5.1 to 5.3 it was easier to point out that these corollaries are indeed consequences of Theorem 4.2. Further, we note that explicit versions of the corollaries can be easily derived from the explicit variant Theorem 4.2* of Theorem 4.2. Finally, the corollaries can be deduced with better bounds from less general versions of Theorem 4.2, where the polynomials $f$ involved are irreducible; for such versions we refer to Győry (1976, 1998, 2000), Evertse and Győry (2017) and in fact Corollary 5.1 above.
5.3. Reformulation of Corollaries 5.1 (ii), 5.2 and 5.3 in terms of polynomial Diophantine equations over $\mathbb{Z}$
a
Let $K$ be an algebraic number field of degree $n\geq 2$ with ring of integers $\mathcal{O}_{K}$ and discriminant $D_{K}$. Let $\{1,\omega_{2},\ldots,\omega_{n}\}$ be an integral basis of $K$. For $\alpha\in\mathcal{O}_{K}$ with
$\alpha=x_{1}+x_{2}\omega_{2}+\cdots+x_{n}\omega_{n},\hskip 8.53581ptx_{1},x_{2},\ldots,x_{n}\in\mathbb{Z},$
its discriminant
(5.3) $D(\alpha)=D(x_{2}\omega_{2}+\cdots+x_{n}\omega_{n})$
can be regarded as a decomposable form of degree $n(n-1)$ in $x_{2},\ldots,x_{n}$ with coefficients in $\mathbb{Z}$, i.e., it is a product of $n(n-1)$ linear forms in $x_{1},\ldots,x_{n}$ with algebraic coefficients. The form $D(x_{2}\omega_{2}+\cdots+x_{n}\omega_{n})$, which was introduced by Kronecker (1882), is called discriminant form, while, for $D\neq 0$, the equation
(5.4) $D(x_{2}\omega_{2}+\cdots+x_{n}\omega_{n})=D\text{ in }x_{2},\ldots,x_{n}\in\mathbb{Z}$
is called a discriminant form equation.
Clearly, Corollary 5.1, (ii) is equivalent to the following
Corollary 5.4 (of Theorem 4.2).
For given $D\neq 0$, the discriminant form equation (5.4) has only finitely many solutions and they can be effectively determined.
The following important fact is due to Hensel (1908):
to the integral basis $\{1,\omega_{2},\ldots,\omega_{n}\}$ of $K$ there corresponds a decomposable form $I(X_{2},\ldots,X_{n})$ of degree $n(n-1)/2$ in $n-1$ variables with coefficients in $\mathbb{Z}$ such that for $\alpha\in\mathcal{O}_{K}$
(5.5) $I(\alpha)=|I(x_{2},\ldots,x_{n})|\text{ {if} }\alpha=x_{1}+x_{2}\omega_{2}+\cdots+x_{n}\omega_{n}$
[…]
Here $I(X_{2},\ldots,X_{n})$ is called an index form, and for given non-zero $I\in\mathbb{Z}$,
(5.6) $I(x_{2},\ldots,x_{n})=\pm I\text{ {in} }x_{2},\ldots,x_{n}\in\mathbb{Z}$
an index form equation.
We note that the equations (5.4) and (5.6) are related by (5.2a).
In view of (5.5), the finiteness assertion of Corollary 5.2 for $\mathcal{O}_{K}$ is equivalent to the following.
Corollary 5.5 (of Theorem 4.2).
For given $I\in\mathbb{Z}\setminus\{0\}$, the index form equation (5.6) has only finitely many solutions, and they can be effectively determined.
In particular, for $I=1$, we get the following equivalent formulation of Corollary 5.3 for $\mathcal{O}_{K}$.
Corollary 5.6 (of Theorem 4.2).
The index form equation
(5.7) $I(x_{2},\ldots,x_{n})=\pm 1\text{ {in} }x_{2},\ldots,x_{n}\in\mathbb{Z}$
has only finitely many solutions, and they can be effectively determined.
The best known upper bound for the solutions of (5.7) is
(5.8) $\max_{2\leq i\leq n}|x_{i}|<\exp\{10^{n^{2}}(|D_{K}|(\log|D_{K}|)^{n})^{n-1}\}$
which is due to Evertse and Győry (2017).
Corollaries 5.5 and 5.6 were proved in Győry (1976) with an explicit upper bound for the sizes of the solutions, not only for equations (5.6) and (5.7) but also for index form equations related to indices with respect to arbitrary orders $\mathcal{O}$ of $K$; see also Győry (2000) and Evertse and Győry (2017).
6. Algorithmic resolution of index form equations, application to (multiply) monogenic number fields
As above, $K$ will denote a number field of degree $n\geq 3$ with ring of integers $\mathcal{O}_{K}$ and discriminant $D_{K}$. For an index form $I(X_{2},\ldots,X_{n})$ associated with an integral basis $\{1,\omega_{2},\ldots,\omega_{n}\}$ of $K$, consider again the above index form equation (5.7).
The exponential bound (5.8) for the solutions of (5.7) is too large for practical use. In the 1990’s, there were new breakthroughs, leading to the complete resolution of certain index form equations. In fact, efficient methods were elaborated for solving equation (5.7) when $|D_{K}|$ is not too large, and the degree $n$ of $K$ is $\leq 6$. Further, (5.7) was solved for many special higher degree number fields $K$ up to about degree $15$ and for some relative extensions of degree $\leq 4$.
6.1. The case $\mathbf{n=3}$ and $\mathbf{4}$. Approach via Thue equations of degree $3$ and $4$.
a
In some cases, the solution of index form equations such as (5.7) can be reduced to one or more Thue equations, i.e., equations of the form $F(x,y)=m$ in $x,y\in\mathbb{Z}$, where $F$ is a binary form with integer coefficients, and $m$ is a non-zero integer. Such an equation is called cubic if $F$ has degree $3$ and quartic if $F$ has degree $4$. For such equations there are efficient algorithms for solving the equations, provided that $H(F)$ and $|m|$ are not too large; see e.g. the books Smart (1998) or Gaál (2019).
For $\mathbf{n=3}$, Gaál and Schulte (1989) reduced (5.7) to a cubic Thue equation. Then, using an efficient algorithm for solving cubic Thue equations, they determined all the solutions of cubic index form equations (5.7), provided that $|D_{K}|$ is not too large.
For $\mathbf{n=4}$, Gaál, Pethő and Pohst (1993, 1996) reduced equation (5.7) to one cubic and some quartic Thue equations. Then, by means of efficient algorithms for solving such Thue equations, they computed all the solutions of equation (5.7) for quartic number fields with not too large discriminant. They obtained several very interesting related results as well.
6.2. Refined version of the general approach combined with reduction and enumeration algorithms
a
For $\mathbf{n\geq 5}$, the approach via Thue equations does not work. In general, for $\mathbf{n=5}$ and $\mathbf{6}$ a refined version of the general approach involving unit equations is needed. Since by (5.5), (5.2a) and (5.1) we have
(5.7)$\Leftrightarrow D_{K/\mathbb{Q}}(\alpha)=D_{K}\Leftrightarrow D(f_{\alpha})=D_{K}\ \text{in\ }\alpha\in\mathcal{O}_{K}$
where $f_{\alpha}\in\mathbb{Z}[X]$ is the minimal polynomial of $\alpha$, in case of concrete equations (5.7) the basic idea of the proof of Theorem 4.2 for irreducible $f_{\alpha}$’s must be combined with some reduction and enumeration algorithms.
The refined version of the general method in the irreducible case consists of the following steps:
  1. (1)
    Reduction to unit equations but in considerably smaller subfields of the normal closure $G$ of $K$. Then the number $r$ of unknown exponents $a_{ijk,l}(1\leq l\leq r)$ in the unit equation (4.5) with $\varepsilon_{ijk}=\zeta_{ijk}\rho_{1}^{a_{ijk,1}}\cdots\rho_{r}^{a_{ijk,r}}$ is much smaller, at most $n(n-1)/2-1$ instead of $r\leq n!-1$; cf. Győry (1998, 2000). Then, in concrete cases one can bound the exponents $|a_{ijk,l}|$ by Baker’s method.
  2. (2)
    The bounds in concrete cases are still too large. Hence a reduction algorithm is needed, reducing the Baker’s bound for $|a_{ijk,l}|$ in several steps if necessary by refined versions of the $L^{3}$-algorithm; cf. de Weger (1989), Wildanger (1997) and Gaál and Pohst (1996).
  3. (3)
    The last step is to apply an enumeration algorithm, determining the small solutions under the reduced bound; cf. Wildanger (1997, 2000), Gaál and Győry (1999) and Bilu, Gaál and Győry (2004).
Combining the refined version of the general method with reduction and enumeration algorithms, for $n=5,6$ and for not too large $|D_{K}|$, Gaál and Győry (1999), resp. Bilu, Gaál and Győry (2004) gave algorithms for determining all power integral bases and hence checking the monogenicity and the multiplicity of the monogenicity of $K$.
We note that the use of the refined version of the general approach is particularly important in the application of the enumeration algorithm.
To perform computations, algebraic number theory packages, a computer algebra system and in some cases a supercomputer were needed.
6.3. Examples: resolutions of index form equations of the form (5.7) for $\mathbf{n=3,4,5,6}$ in the most difficult case
a
In the examples below, the authors resolved concrete index form equations of the form (5.7) for $n=3,4,5,6$. The involved number fields $K$ of degree $n$ are given by irreducible monic polynomials $f(X)\in\mathbb{Z}[X]$, a zero of which generates the corresponding $K$ over $\mathbb{Q}$. In each case all power integral bases in $K$, and therefore the multiplicity of the monogenicity of $K$, denoted by $mm(K)$, are computed by the method outlined above. For the lists of the power integral bases, we refer to the original papers and to Evertse and Győry (2017) and Gaál (2019).
  • $\mathbf{n=3}$, $f(X)=X^{3}-X^{2}-2X+1$, $mm(K)=9$ (Gaál and Schulte, 1989);
  • $\mathbf{n=4}$, $f(X)=X^{4}-4X^{2}-X+1$, $mm(K)=17$ (Gaál, Pethő and Pohst, 1990’s);
  • $\mathbf{n=5}$, $f(X)=X^{5}-5X^{3}+X^{2}+3X-1$, $mm(K)=39$ (Gaál and Győry, 1999);
  • $\mathbf{n=6}$, $f(X)=X^{6}-5X^{5}+2X^{4}+18X^{3}-11X^{2}-19X+1$, $mm(K)=45$ (Bilu, Gaál, and Győry, 2004);
We note that from the point of view of computation, the above examples belong to the most difficult cases, $K$ being in each case totally real with Galois group $S_{n}$. In these cases the number of exponents in the unit equations involved is the largest possible.
For $n\geq 7$, the above mentioned algorithms do not work in general. Then the number of fundamental units, $\rho_{1},\ldots,\rho_{r}$ involved can be too large to use the enumeration algorithm.
Problem 1.
For $\mathbf{n=7}$, give a practical algorithm for solving equation (5.7) in case of any number field $K$ of degree $7$ with not too large discriminant.
7. Power integral bases and canonical number systems in number fields
Number systems and their generalizations have been intensively studied for a long time. Here we present an important generalization for the number field case, point out its close connection with power integral bases and formulate an application of the above Corollary 5.3 to this generalization.
Let $K$ be an algebraic number field with ring of integers $\mathcal{O}_{K}$, and let $\alpha\in\mathcal{O}_{K}$ with $|N_{K/\mathbb{Q}}(\alpha)|\geq 2$. Then $\{\alpha,\mathcal{N}(\alpha)\}$ with
$\mathcal{N}(\alpha)=\{0,1,\ldots,|N_{K/\mathbb{Q}}(\alpha)-1\}$
is called a canonical number system, in short CNS, in $\mathcal{O}_{K}$, if every non-zero element of $\mathcal{O}_{K}$ has a unique representation of the form
$a_{0}+a_{1}\alpha+\cdots+a_{k}\alpha^{k}\text{ with }a_{i}\in\mathcal{N}(\alpha)\text{ for }i=0,\ldots,k,\,a_{k}\neq 0$
Then $\alpha$ is called the base and $\mathcal{N}(\alpha)$ the set of digits of the number system. This concept is a generalization of the radix representation considered in $\mathbb{Z}$.
Kovács (1981) proved the following fundamental result.
Theorem 7.1 (Kovács, 1981).
In $\mathcal{O}_{K}$ there exists a canonical number system if and only if $\mathcal{O}_{K}$ has a power integral basis.
Together with the above Corollary 5.3 this implies that it is effectively decidable whether there exists a CNS in $\mathcal{O}_{K}$. Corollary 5.3 provides even a general algorithm to determine all power integral bases in $\mathcal{O}_{K}$. Using this, Kovács and Pethő (1991) characterized the bases of all CNS’s of $\mathcal{O}_{K}$. As a consequence they showed as follows.
Theorem 7.2 (Kovács and Pethő, 1991).
Up to $\mathbb{Z}$-equivalence, there are only finitely many CNS’s in $\mathcal{O}_{K}$, and all of them can be effectively determined.
In fact, they extended their result to any order of $K$ as well.
We note that Brunotte (2001) considerably improved the procedure of Kovács and Pethő (1991) and gave an efficient algorithm for finding all such CNS’s, provided that one has an efficient algorithm for determining all power integral bases in $\mathcal{O}_{K}$. As was seen in Section 6, such an efficient algorithm is known for number fields $K$ of degree at most $6$ if their discriminants are not too large in absolute value.
For surveys on CNS’s, we refer to Brunotte (2001), Pethő (2004), Brunotte, Huszti and Pethő (2006), Evertse, Győry, Pethő and Thuswaldner (2019) and the references given there.
8. Generalizations and further applications
For a detailed, comprehensive treatment of the reduction theory of integral polynomials with given non-zero discriminant, including Theorems 4.1, 4.2, 4.4 and their generalizations and applications, we refer to the monograph Evertse and Győry (2017) and the references given there. For the algorithmic aspect one can consult also the books Smart (1998) and Gaál (2019), and for general results over integral domains of characteristic 00 that are finitely generated over $\mathbb{Z}$, the book Evertse and Győry (2022).
Below we present, resp. mention some of the most important generalizations and applications.
8.1. Generalizations
a
Theorems 4.1 and 4.4 were proved in Birch and Merriman (1972) resp. in Evertse and Győry (1991) in more general forms, over number fields, more precisely over the $S$-integers of a number field. Theorem 4.2, (ii) was extended to this more general situation in Győry (1978a, 1981).
We present a generalization of Corollary 5.3 in the relative case. Let $L$ be a number field, $K/L$ a finite relative extension, $\mathcal{O}_{L}$ and $\mathcal{O}_{K}$ the rings of integers of $L$ and $K$. Then $K/L$ is called monogenic if $\mathcal{O}_{K}=\mathcal{O}_{L}[\alpha]$ for some $\alpha\in\mathcal{O}_{K}$, and $\{1,\alpha,\ldots,\alpha^{n-1}\}$ is a relative power integral basis of $K$ over $L$, where $n=[K:L]$, i.e., an $\mathcal{O}_{L}$-module basis of $\mathcal{O}_{K}$. We say that $\alpha,\beta\in\mathcal{O}_{K}$ are $\mathcal{O}_{L}$-equivalent if $\beta=a+\varepsilon\alpha$ for some $a\in\mathcal{O}_{L}$ and unit $\varepsilon$ in $\mathcal{O}_{L}$. If $\alpha$ is a generator of $\mathcal{O}_{K}$ over $\mathcal{O}_{L}$, i.e., $\mathcal{O}_{K}=\mathcal{O}_{L}[\alpha]$ then so is every $\beta$ $\mathcal{O}_{K}$-equivalent to $\alpha$.
Theorem 8.1 (Győry, 1978a).
There are only finitely many $\mathcal{O}_{L}$-equivalence classes of $\alpha\in\mathcal{O}_{K}$ with $\mathcal{O}_{K}=\mathcal{O}_{L}[\alpha]$, and a full set of representatives of such $\alpha$ can be, at least in principle, effectively determined.
For the best known version with explicit bounds, see Corollary 8.4.13 in Evertse and Győry (2017).
We now present two general finiteness theorems where the ground ring is an integrally closed integral domain $A$ of characteristic 00 that is finitely generated over $\mathbb{Z}$ as a $\mathbb{Z}$-algebra, i.e., $A=\mathbb{Z}[z_{1},\ldots,z_{r}]$, where we allow some of the $z_{i}$ to be transcendental. The unit group of $A$ is denoted by $A^{*}$.
We say that the monic polynomials $f,g\in A[X]$ are $A$-equivalent if $g(X)=f(X+a)$ with some $a\in A$. Then $f$ and $g$ have the same discriminant.
Theorem 8.2 (Győry, 1982).
Let $G$ be a finite extension of the quotient field of $A$. Up to $A$-equivalence, there are only finitely many monic $f(X)$ in $A[X]$ with given non-zero discriminant $D$ having all their zeros in $G$.
This was made effective by Győry (1984) in a special case, and in full generality by Evertse and Győry (2017), provided that $A,G$ and $D$ are given effectively in the sense defined in Evertse and Győry (2017, 2022).
Theorem 8.3 (Evertse and Győry, 2017, 2022).
Up to $A$-equivalence, there are only finitely many monic $f$ in $A[X]$ with $D(f)=D$, and if $A,G,D$ are effectively given, all these $f$ can be effectively determined.
Problem 2.
Are these statements true without fixing the splitting field $G$?
  • Theorem 4.4 was later generalized for decomposable forms in more than two variables in Evertse and Győry (1992) and Győry (1994).
  • Theorem 4.2, (ii) and Corollaries 5.4, 5.6 were extended to the case when $D$ resp. $I$ is replaced by $p_{1}^{z_{1}},\ldots,p_{s}^{z_{s}}$, where $p_{1},\ldots,p_{s}$ are fixed primes and $z_{1},\ldots,z_{s}$ unknown non-negative integers; see Győry (1978b, 1981), Trelina (1977a, 1977b), Győry and Papp (1977). These results yielded e.g. explicit lower bound for the greatest prime factor of discriminant and index of an integer of a number field. For generalizations for the number field case, see Győry (1980a).
  • Corollary 5.4 was generalized for more general discriminant form equations
    $D(x_{2}\omega_{2}+\cdots+x_{k}\omega_{k})=D\text{ in }x_{2},\ldots,x_{k}\in\mathbb{Z}$
    where $1,\omega_{2},\ldots,\omega_{k}$ are $\mathbb{Q}$-linearly independent elements of a number field $K$ of degree $n$ with $k\leq n$; see e.g. Győry (1976) and Győry (2000) with a further generalization for the number field case.
  • Corollary 5.4 on discriminant form equations was generalized for more general decomposable form equations of the form
    (8.1) $F(x_{1},\ldots,x_{m})=F\text{ in }x_{1},\ldots,x_{m}\in\mathbb{Z},$
    where $F\in\mathbb{Z}\setminus\{0\}$ and $F(X_{1},\ldots,X_{m})$ is a decomposable form with coefficients in $\mathbb{Z}$ which factorizes into linear factors over $\overline{\mathbb{Q}}$ such that these factors form a so-called triangularly connected system (i.e. (8.1) can be reduced to a connected system of three terms unit equations); see Győry and Papp (1978) and, more generally, Győry (1998).
  • Corollary 5.4 was generalized for the ‘inhomogeneous’ case by Gaál (1986).
  • Analogous results were established over function fields by Győry (1984, 2000); Gaál (1988), Mason (1988), Shlapentokh (1996).
Several results of the theory have been extended to the case of étale algebras in Evertse and Győry (2017, 2022).
  • Let $K$ be a number field with ring of integers $\mathcal{O}_{K}$, and $D\neq 0$ an integer. It is a special case of Corollary 5.1 that, up to $\mathbb{Z}$-equivalence, the equation
    (8.2) $D(\alpha)=D\text{ in }\alpha\in\mathcal{O}_{K}$
    has only finitely many solutions, and that all of these can be effectively determined.
    Let $A=\mathbb{Z}[z_{1},\ldots,z_{r}]$ be an integral domain of characteristic 00 with algebraic or transcendental generators $z_{1},\ldots,z_{r}$, $L$ its quotient field, and $\Omega$ a finite étale $L$-algebra (i.e. a direct product of finite extensions $K_{1},\ldots,K_{t}$ of $L$). Denote by $A_{\Omega}$ the integral closure of $A$ in $\Omega$. The discriminant of $\alpha\in A_{\Omega}$ over $L$ with $\Omega=L[\alpha]$ is given by $D_{L}(\alpha):=D(f_{\alpha})$, where $f_{\alpha}$ is the monic minimal polynomial of $\alpha$ over $L$.
    Let $\mathcal{O}$ be an $A$-order of $\Omega$, i.e. an $A$-subalgebra of $A_{\Omega}$ which spans $\Omega$ as an $L$-vector space.
    Slightly inconsistently with earlier given definitions, we say that $\alpha,\beta\in\mathcal{O}$ are $A$-equivalent if $\beta-\alpha\in A$. One verifies that if $\alpha,\beta\in\mathcal{O}$ are $A$-equivalent then $f_{\alpha}$, $f_{\beta}$ are $A$-equivalent, and thus, $D_{L}(\beta)=D_{L}(\alpha)$.
    Let $D$ be a non-zero element of $L$. Consider the following generalization of equation (8.2):
    (8.3) $D_{L}(\alpha)=D\text{ in }\alpha\in\mathcal{O}.$
    For an integral domain $B$, denote by $B^{+}$ the additive group of $B$.
    Theorem 8.4 (Evertse and Győry, 2022).
    If
    (8.4) $(\mathcal{O}\cap L)^{+}/A^{+}\text{{ is finite,}}$,
    then the set of $\alpha\in\mathcal{O}$ with (8.3) is a union of finitely many $A$-equivalence classes. Moreover, if $A,\Omega,\mathcal{O}$ and $D$ are given effectively in a well-defined way, one can determine a set consisting of precisely one element from each of these classes.
    The condition (8.4) is necessary and decidable.
    For $A=\mathbb{Z}$, $L=\mathbb{Q}$, $\Omega=$ number field $K$, $\mathcal{O}=\mathcal{O}_{K}$, Theorem 8.4 gives the above theorem concerning equation (8.3).
8.2. Further applications
a
Theorems 4.1, 4.2 and 4.4 as well as their various versions and generalizations led to many applications. Some of them were treated in Sections 4 to 7. Below we briefly present some others in their simplest form. For further applications, we refer to the survey papers Győry (1980a, 2000, 2006), the books Győry (1980b), Smart (1998), Evertse and Győry (2017, 2022), Gaál (2019), and the references given there. Applications to classical Diophantine equations.
  • Let $f\in\mathbb{Z}[X]$ be a monic polynomial of degree $n\geq 3$ with discriminant $D(f)\neq 0$, and $m\geq 2$ an integer. Consider the solutions $x,y\in\mathbb{Z}$ of the equation
    (8.5) $f(x)=y^{m}.$
    Using various variants of Theorem 4.2 (ii), Trelina (1985) and, for $n=3,m=2$, Pintér (1995) gave effective upper bounds for $|y|$ that depend on $m,n$ and $|D(f)|$, but not on the height of $f$. We recall that the height of $f$ can be arbitrarily large with respect to $|D(f)|$. Furthermore, Győry and Pintér (2008) showed that for each solution $x,y$ of (8.5) with ${\rm gcd}(y,D(f))=1$, $|y|^{m}$ can be effectively bounded in terms of the radical of $D(f)$, i.e. the product of the distinct prime factors of $D(f)$. Brindza, Evertse and Győry (1991), Haristoy (2003) and Győry and Pintér (2008) gave upper bounds even for $m$ that depend only on $n$ and $|D(f)|$.
  • Let $F\in\mathbb{Z}[X,Y]$ be an irreducible binary form of degree $n\geq 3$ and discriminant $D$, let $p_{1},\ldots,p_{s}\,\,(s\geq 0)$ be distinct primes at most $P$, and let $m$ be a positive integer coprime with $p_{1},\ldots,p_{s}$. There are several upper bounds for the number of solutions $x,y$ of the Thue equation
    (8.6) $F(x,y)=m,$
    the Thue inequality
    (8.7) $0<|F(x,y)|\leq m$
    and the Thue–Mahler equation
    (8.8) $F(x,y)=mp_{1}^{z_{1}}\cdots p_{s}^{z_{s}}\text{, with }(x,y)=1,$
    where $z_{1},\ldots,z_{2}$ are also unknown non-negative integers.
    Using a quantitative version of Theorem 4.4, i.e. the general effective result of Evertse and Győry (1991a) on binary forms of given degree and given discriminant, previously obtained upper bounds for the number of solutions of these equations were substantially improved under the assumptions that $n,D,m,s$ and $P$ satisfy some additional conditions. Such improved upper bounds were derived in Stewart (1991) for (8.8) with ${\rm gcd}(x,y)=1$ when $m>C_{1}$, in Brindza (1996) for (8.6) with ${\rm gcd}(x,y)=1$ when $m>C_{2}$, and in Thunder (1995) for (8.7) when $m>C_{3}$, where $C_{1},C_{2},C_{3}$ are effectively computable numbers such that $C_{1}$ depends on $n,|D|,P,s$ and $C_{2},C_{3}$ on $n$ and $|D|$. Further, Evertse and Győry (1991b) showed that if $|D|>C_{4}$, then the number of coprime solutions of (8.7) is at most $6n$ if $n>400$, and by Győry (2001) it is at most $628n+$ 6 if $|D|>C_{5}$ and $3\leq n\leq 400$. For $m=1$ and $|D|>C_{6}$, this was later improved by Akhtari (2012) to $211n-$ 2. Here $C_{4},C_{5},C_{6}$ are effectively computable numbers such that $C_{4},C_{5}$ depend on $m$ and $n$, and $C_{6}$ on $n$. Together with a quantitative version of Theorem 4.4, i.e. with the result of Evertse and Győry (1991a), these imply that for given $n\geq 3$ and $m\geq 1$, there are only finitely many $GL_{2}(\mathbb{Z})$-equivalence classes of irreducible binary forms $F\in\mathbb{Z}[X,Y]$ of degree $n$ for which the number of coprime solutions of (8.7) exceeds $628n+$ 6 or $211n-$ 2 if $m=1$.
  • The quantitative version of Theorem 4.4, proved in Evertse and Győry (1991a) was also applied in Evertse (1993) to bound the number of solutions of some resultant inequalities, and in Ribenboim (2006) to binary forms with given discriminant, having additional conditions on the coefficients. We remark that using the improved and completely explicit version Theorem 4.4* of Evertse and Győry (2017), the above quoted applications can be made more precise.
Further applications of Theorems 4.2 and 4.4:
  • Some applications of Theorem 4.2 were given to the reducibility of a general class of polynomials of the form $g(f(X))$ where $f,g$ are monic polynomials, $g(X)$ is irreducible with CM splitting field. For given prime $p$, there are up to $\mathbb{Z}$-equivalence only finitely many $f\in\mathbb{Z}[X]$ of degree $p$ with distinct real zeros for which $g(f(X))$ is reducible; see Győry (1976, 1982).
  • For an application of an earlier version of Corollary 5.1 (ii) to integral valued polynomials over the set of algebraic integers of bounded degree, see Peruginelli (2014).
  • For an application of Corollary 5.1 to so-called binomially equivalent numbers, see Yingst (2006).
  • Let $K/L$ be a field extension of degree $n\geq 2$, and $\mathcal{O}_{K},\mathcal{O}_{L}$ the rings of integers of $K$ resp. $L$. Pleasants (1974) gave an explicit formula which enables one to compute a positive integer $m(\mathcal{O}_{K},\mathcal{O}_{L})$ such that if $r(\mathcal{O}_{K},\mathcal{O}_{L})$ denotes the minimal number of generators of $\mathcal{O}_{K}$ as $\mathcal{O}_{L}$-algebra then
    (8.9) $m(\mathcal{O}_{K},\mathcal{O}_{L})\leq r(\mathcal{O}_{K},\mathcal{O}_{L})\leq\max\{m(\mathcal{O}_{K},\mathcal{O}_{L}),2\}.$
    Pleasants proved that if $L=\mathbb{Q}$, there are number fields $K$ of arbitrarily large degree over $\mathbb{Q}$ such that $m(\mathcal{O}_{K},\mathbb{Z})=1$ and $\mathcal{O}_{K}$ is not monogenic. Consequently, his theorem does not make it possible to decide whether the ring of integers of a number field is monogenic. Together with Pleasants’ result, our Theorem 8.1 above gives the following
    Corollary of Theorem 8.1 (and Pleasants (1974)).
    There is an algorithm for determining the least number of elements of $\mathcal{O}_{K}$ that generate $\mathcal{O}_{K}$ as an $\mathcal{O}_{L}$-algebra.
    Chapter 11 of Evertse and Győry (2017) considers more generally $\mathcal{O}_{S}$-orders of finite étale $L$-algebras, and gives a method to determine a system of $\mathcal{O}_{S}$-algebra generators of minimal cardinality of such an order. This was basically work of Kravchenko, Mazur and Petrenko (2012), worked out in more detail in a special case.
Applications of Theorems 4.4 and 4.4*:
  • In Evertse and Győry (1991a), effective upper bounds were given for the minimal non-zero absolute value of binary forms at integral points.
  • An effective upper bound was derived for the height of appropriate representatives of $GL_{2}(\mathbb{Z})$-equivalence classes of algebraic numbers with given discriminant.
  • In Evertse and Győry (2017), as a consequence of Theorem 4.4*, we derived for any separable polynomial $f\in\mathbb{Z}[X]$ of degree $n\geq 4$ an improvement of the previous bounds for the minimal root distance of $f$.
  • Evertse (2023) applied a quantitative version of Theorem 4.4 to so-called rational monogenizations of orders of a number field. For further details we refer to Appendix B.
  • Theorem 4.4 can be generalized to a statement for polynomials in $\mathcal{O}_{S}[X]$, where $S$ is a finite set of places on a number field $L$ containing the infinite ones, and $\mathcal{O}_{S}$ is the ring of $S$-integers of $L$. Using a quantitative version of such a generalization of Evertse and Győry (1991a), von Känel (2011, 2014) proved an effective version of Shafarevich’ conjecture/Faltings’ theorem for hyperelliptic curves.
  • Evertse and Győry (2017), following von Känel’s proof, gave an improved and completely explicit version of von Känel’s theorem.
  • Szpiro and Tucker (2008) established an analogue for self-maps of the projective line over a number field of Shafarevich’ conjecture on the reduction of algebraic curves modulo primes, cf. Shafarevich (1963). Petsche (2012) proved an analogue of Shafarevich’ conjecture for families of critically separable rational maps over number fields.
  • Evertse and Győry (1992) applied their effective finiteness theorem on decomposable forms of given discriminant to decomposable form equations. Their result was used by Stout (2014) to prove that for a given number field $K$, finite set of places $S$ of $K$ and rational morphism $\Phi:\,\mathbb{P}^{n}\rightarrow\mathbb{P}^{n}$ defined over $K$, there are only finitely many twists of $\Phi$ defined over $K$ which have good reduction at all places outside $S$. This answered a question of Silverman in the affirmative.
Appendix: Related topics
We discuss some topics related to monogenic number fields and monogenic orders and generalizations thereof that do not strictly belong to the reduction theory of integral polynomials.
A. Multiply monogenic orders
In this section, we give upper bounds for the multiplicity of monogenicity of orders of number fields.
Let $K$ be a number field with ring of integers $\mathcal{O}_{K}$, and $\mathcal{O}$ an arbitrary order of $K$, i.e., a subring of $\mathcal{O}_{K}$ with quotient field $K$. It follows from Corollary 5.3 above that up to $\mathbb{Z}$-equivalence, there are only finitely many $\alpha$ in $\mathcal{O}$ with $\mathcal{O}=\mathbb{Z}[\alpha]$. The order $\mathcal{O}$ is said to be $k$-times monogenic/precisely $k$ times monogenic/at most $k$ times monogenic if there are at least/precisely/at most $k$ pairwise $\mathbb{Z}$-inequivalent such generators $\alpha$ of $\mathcal{O}$ over $\mathbb{Z}$.
It is easy to see that every order of a quadratic number field is precisely one time monogenic.
For $n\geq 3$, Evertse and Győry (1985) proved that every order of a number field $K$ of degree $n$ is at most $(3\times 7^{2g})^{n-2}$-times monogenic, where $g$ denotes the degree of the normal closure of $K$. Note that $n\leq g\leq n!$. This was the first uniform bound of this type.
For fixed $n\geq 3$, we denote by $M(n)$ the smallest integer $k$ such that for every number field $K$ of degree $n$ and every order $\mathcal{O}$ of $K$, the order $\mathcal{O}$ is at most $k$ times monogenic. The above result of Evertse and Győry implies that for $n\geq 3$, the quantity $M(n)$ is finite and in fact, $M(n)\leq(3\times 7^{2n!})^{n-2}$.
The problem of estimating $M(3)$ can be reduced via index form equations to estimating the number of integer solutions of a Thue equation $|F(x,y)|=1$ with $F$ an integral cubic binary form. Bennett (2001) proved that such an equation has up to sign at most $10$ solutions. This gives the following.
Theorem A.1 (Bennett, 2001).
We have $M(3)\leq 10$.
Then, for $n\geq 4$, Evertse (2011) improved the bound of Evertse and Győry (1985).
Theorem A.2 (Evertse, 2011).
For $n\geq 4$, $M(n)\leq 2^{4(n+5)(n-2)}$ holds.
The main tool in the proof of the general bound is the following result of Beukers and Schlickewei (1996): Let $a,b$ be non-zero complex numbers and $\Gamma$ a multiplicative subgroup of $\mathbb{C}^{*}$ of rank $r$. Then the equation $ax+by=1$ has at most $2^{16r+8}$ solutions in $x,y\in\Gamma$.
In the case of quartic number fields, Bhargava (2022) substantially improved Evertse’s bound by proving the following theorem.
Theorem A.3 (Bhargava, 2022).
We have $M(4)\leq 2760$ (and $M(4)\leq 182$ if $|D(\mathcal{O})|$ is sufficiently large).
Bhargava proved his theorem via a parametrization of quartic rings and their cubic resolvent rings, and utilized Akhtari’s recent upper bound (see the Appendix of Bhargava (2022)) for the number of solutions of quartic Thue equations.
Akhtari (2022) gave another, more direct proof for Theorem A.3, following the approach of Gaál, Pethő and Pohst (1996) (which in fact is going into the same direction as Bhargava’s approach but is less general), and combining this with her own upper bound for the number of solutions of quartic Thue equations.
Theorem A.2 is probably far from best possible in terms of $n$. We pose the following problem:
Problem 3 (Győry, 2000).
Do there exist absolute constants $c_{1},c_{2}$ such that $M(n)<c_{1}n^{c_{2}}$ for all $n\geq 4$
We now fix a number field $K$ of degree $\geq 3$, and consider only orders of $K$. As it turned out, most orders of $K$ have only small multiplicity of monogenicity, bounded above independently even of the degree of $K$. In 2013, we proved the following result with Bérczes:
Theorem A.4 (Bérczes, Evertse and Győry, 2013).
Let $K$ be an algebraic number field of degree $\geq 3$. Then $K$ has only finitely many orders that are three times monogenic.
To see that the bound $3$ is optimal, let $K$ be a non-CM number field of degree $\geq 3$. Then the ring of integers of $K$ has infinitely many units $\varepsilon$ with $K=\mathbb{Q}(\varepsilon)$. For every of these $\varepsilon$ we obtain a two times monogenic order $\mathbb{Z}[\varepsilon]=\mathbb{Z}[\varepsilon^{-1}]$ of $K$.
Theorem A.4 is proved by means of a reduction to unit equations in more than two unknowns, and a use of ineffective finiteness theorems for these equations. So Theorems A.4 is ineffective, in the sense that its proof does not allow to determine the exceptional orders.
Problem 4.
Make Theorem A.4 effective.
This seems to be completely out of reach. At present, it is not known how to make the results on unit equations in more than two unknowns effective.
B. Generalizations to rationally monogenic orders
The results in Appendix A deal with monogenic orders, i.e., orders of the shape $\mathbb{Z}[\alpha]$ where $\alpha$ is an algebraic integer. There are generalizations of such orders, denoted $\mathbb{Z}_{\alpha}$, called rationally monogenic orders, attached to non-integral algebraic numbers $\alpha$. Further, the theorems stated in Appendix A have analogues for rationally monogenic orders. Before we can define rationally monogenic orders and state our theorems, we briefly go into some history and introduce the necessary terminology.
While in the results for monogenic orders, $\mathbb{Z}$-equivalence of algebraic integers plays an important role, for rationally monogenic orders we have to deal with $GL_{2}(\mathbb{Z})$-equivalence of algebraic numbers. Two algebraic numbers $\alpha,\beta$ are called $GL_{2}(\mathbb{Z})$-equivalent if $\beta=\frac{a\alpha+b}{c\alpha+d}$ for some matrix $\big{(}\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big{)}\in GL_{2}(\mathbb{Z})$. The correspondence with polynomials is as follows. Denote by $f_{\alpha}$ the primitive minimal polynomial of an algebraic number $\alpha$, i.e.,
(B.1) $f_{\alpha}=a_{0}X^{n}+\cdots+a_{n}=a_{0}(X-\alpha^{(1)})\cdots(X-\alpha^{(n)})\in\mathbb{Z}[X]$
where $a_{0}>0$, ${\rm gcd}(a_{0},\ldots,a_{n})=1$ and $\alpha^{(1)}=\alpha,,\ldots,\alpha^{(n)}$ are the conjugates of $\alpha$. Then if $\alpha,\beta$ are $GL_{2}(\mathbb{Z})$-equivalent then so are $f_{\alpha}$, $f_{\beta}$ while conversely, if $f_{\alpha},f_{\beta}$ are $GL_{2}(\mathbb{Z})$-equivalent, then $\alpha$ is $GL_{2}(\mathbb{Z})$-equivalent to a conjugate of $\beta$.
Let $\alpha$ be a non-zero, not necessarily integral algebraic number of degree $n\geq 3$, and $f_{\alpha}$ its primitive minimal polynomial given by (B.1). Define $\mathbb{Z}_{\alpha}$ to be the $\mathbb{Z}$-module with basis
$1,\ \omega_{2}:=a_{0}\alpha,\ \omega_{3}:=a_{0}\alpha^{2}+a_{1}\alpha,\ldots,\omega_{n}:=a_{0}\alpha^{n-1}+a_{1}\alpha^{n-2}+\cdots+a_{n-2}\alpha.$
This $\mathbb{Z}$-module was introduced by Birch and Merriman (1972), who observed that it is contained in the ring of integers of $\mathbb{Q}(\alpha)$, and that its discriminant is equal to the discriminant $D(f_{\alpha})$ of $f_{\alpha}$. Nakagawa (1989) showed that it is in fact an order of the field $\mathbb{Q}(\alpha)$, i.e., closed under multiplication. This order was further studied by Simon (2001, 2003) and Del Corso, Dvornicich and Simon (2005). As was very likely known at the time, the order $\mathbb{Z}_{\alpha}$ can be described more concisely as follows. Let $\mathcal{M}_{\alpha}$ be the $\mathbb{Z}$-module generated by $1,\alpha,\ldots,\alpha^{n-1}$. Then
$\mathbb{Z}_{\alpha}=\left\{\xi\in\mathbb{Q}(\alpha):\ \xi\mathcal{M}_{\alpha}\subseteq\mathcal{M}_{\alpha}\right\}.$
In the case that $\alpha$ is an algebraic integer of degree $n$, the powers $\alpha^{i}$ ($i\geq n$) belong to $\mathcal{M}_{\alpha}$, and thus, $\mathbb{Z}_{\alpha}=\mathcal{M}_{\alpha}=\mathbb{Z}[\alpha]$. Further, if $\alpha$, $\beta$ are non-zero $GL_{2}(\mathbb{Z})$-equivalent algebraic numbers, then $\mathbb{Z}_{\alpha}=\mathbb{Z}_{\beta}$. Indeed, let $\beta=\frac{a\alpha+b}{c\alpha+d}$ for some matrix $\big{(}\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big{)}\in GL_{2}(\mathbb{Z})$. Then $\mathcal{M}_{\beta}=(c\alpha+d)^{1-n}\mathcal{M}_{\alpha}$ where $n=\deg\alpha$, and thus, $\mathbb{Z}_{\alpha}=\mathbb{Z}_{\beta}$.
Nowadays, for a non-zero algebraic number $\alpha$ we call $\mathbb{Z}_{\alpha}$ the invariant order or invariant ring of $f_{\alpha}$.
From the effective result of Evertse and Győry (1991a) and the fact that $\mathbb{Z}_{\alpha}$ has the same discriminant as $f_{\alpha}$ the following analogue of Corollary 5.3 can be deduced:
Theorem B.1.
Let $\mathcal{O}$ be an order of a number field $K$ of degree $n$. Denote by $D_{\mathcal{O}}$ the discriminant of $\mathcal{O}$. Every $\alpha$ such that $\mathbb{Z}_{\alpha}=\mathcal{O}$ is $GL_{2}(\mathbb{Z})$-equivalent to some $\beta\in K$ of height $H(\beta)\leq c_{1}(n,|D_{\mathcal{O}}|)$, where $c_{1}(n,|D_{\mathcal{O}}|)$ is effectively computable in terms of $n$ and $|D_{\mathcal{O}}|$.
A consequence is that up to $GL_{2}(\mathbb{Z})$-equivalence there are only finitely many $\alpha\in K$ such that $\mathbb{Z}_{\alpha}=\mathcal{O}$, and that these can be effectively determined.
We call an order $\mathcal{O}$ of a number field $K$ rationally monogenic if there is $\alpha$ such that $\mathcal{O}=\mathbb{Z}_{\alpha}$. Clearly, monogenic orders are rationally monogenic. In Evertse (2023) it was shown that every number field $K$ of degree $\geq 3$ has infinitely many orders that are rationally monogenic but not monogenic. Simon (2001) gave various examples of number fields of degree $\geq 4$ that are not rationally monogenic, i.e., whose rings of integers are not rationally monogenic.
We say that an order $\mathcal{O}$ of a number field $K$ is $k$ times/precisely $k$ times/at most $k$-times rationally monogenic if up to $GL_{2}(\mathbb{Z})$-equivalence there are at least/precisely/at most $k$ numbers $\alpha$ such that $\mathcal{O}=\mathbb{Z}_{\alpha}$. Denote by $RM(n)$ the least number $k$ such that for every number field $K$ of degree $n$ and every order $\mathcal{O}$ of $K$, the order $\mathcal{O}$ is at most $k$ times rationally monogenic.
From work of Delone and Faddeev (1940) it follows that $RM(3)\leq 1$, that is, every order of a cubic number field is at most one time rationally monogenic. From a result of Bérczes, Evertse and Győry (2004) the following analogue of Theorem A.1 can be deduced:
Theorem B.2.
For every $n\geq 4$, $RM(n)$ is finite and in fact, $RM(n)\leq n\times 2^{24n^{3}}$.
Similarly to Theorem A.2 the proof uses the result of Beukers and Schlickewei (1996) mentioned above.
This bound has been improved. The best bounds to date are as follows:
Theorem B.3.
We have
  • (i)
    $RM(4)\leq 40$ (Bhargava (2022));
  • (ii)
    $RM(n)\leq 2^{5n^{2}}$ for $n\geq 5$ (Evertse and Győry (2017)).
The proof of part (ii) is similar to that of Theorem B.2 but with a combinatorial improvement in the argument. The proof of part (i) also uses a parametrization of quartic rings and their cubic resolvent rings.
Recently, the following analogue of Theorem A.4 for rationally monogenic orders was proved:
Theorem B.4 (Evertse, 2023).
  • (i)
    Let $K$ be a number field of degree $4$. Then $K$ has only finitely many three times rationally monogenic orders.
  • (ii)
    Let $K$ be a number field of degree $\geq 5$ such that the normal closure of $K$ is $5$-transitive. Then $K$ has only finitely many two times rationally monogenic orders.
Part (i) is best possible in the sense that there are quartic number fields having infinitely many two times rationally monogenic orders. It is not clear whether the $5$-transitivity condition on the Galois closure of $K$ in part (ii) is necessary; this was just a technical condition needed for the proof.
Similary to Theorem A.4, Theorem B.4 has been proved by means of a reduction to unit equations in more than two unknowns, and a use of ineffective finiteness theorems for such equations. So likewise, Theorem B.4 is ineffective.
We would like to finish with a connection with Hermite equivalence classes, discussed in Subsection 3.2. Let $\alpha$ be an algebraic number of degree $n$. Let $\mathcal{I}_{\alpha}:=\mathbb{Z}_{\alpha}+\alpha\mathbb{Z}_{\alpha}$ be the fractional ideal of $\mathbb{Z}_{\alpha}$ generated by $1$ and $\alpha$. This is known to be invertible, see Simon (2003). It is called also the invariant ideal of $f_{\alpha}$.
Recall that a polynomial in $\mathbb{Z}[X]$ is primitive if its coefficients have greatest common divisor $1$.
Theorem B.5 (BEGyRS, 2023).
Let $f,g\in\mathbb{Z}[X]$ be two primitive, irreducible polynomials. Then the following two assertions are equivalent:
  • (i)
    $f$ and $g$ are Hermite equivalent;
  • (ii)
    $f$ has a root $\alpha$ and $g$ a root $\beta$ such that $\mathbb{Z}_{\alpha}=\mathbb{Z}_{\beta}$ and $\mathcal{I}_{\alpha}$ and $\mathcal{I}_{\beta}$ lie in the same ideal class of $\mathbb{Z}_{\alpha}$.
In the particular case that $f$ and $g$ are monic, we have $\alpha\in\mathbb{Z}[\alpha]=\mathbb{Z}_{\alpha}$ and $\mathcal{I}_{\alpha}=\mathbb{Z}_{\alpha}$ and likewise for $g$ and $\beta$. So two monic, irreducible polynomials $f,g\in\mathbb{Z}[X]$ are Hermite equivalent if and only if $f$ has a root $\alpha$ and $g$ a root $\beta$ such that $\mathbb{Z}[\alpha]=\mathbb{Z}[\beta]$.
Combining Theorems B.5 and B.4 one can deduce the following counterpart of Theorem 3.4. For a number field $K$, let $\mathcal{P}\mathcal{I}(K)$ denote the set of primitive, irreducible polynomials in $\mathbb{Z}[X]$ having a root $\alpha$ such that $K=\mathbb{Q}(\alpha)$.
Theorem B.6 (Evertse, 2023).
  • (i)
    Let $K$ be a quartic number field. Then $\mathcal{P}\mathcal{I}(K)$ has only finitely many Hermite equivalence classes that split into more than two $GL_{2}(\mathbb{Z})$-equivalence classes.
  • (ii)
    Let $K$ be a number field of degree $\geq 5$ whose normal closure is $5$-transitive. Then $\mathcal{P}\mathcal{I}(K)$ has only finitely many Hermite equivalence classes that split into more than one $GL_{2}(\mathbb{Z})$-equivalence class.
Part (ii) was conjectured in BEGyRS (2023), without the $5$-transitivity condition.
In a later, yet to be written extended version of this paper, we will give more details and background about the above mentioned theorems.
C. Monogenicity, class group and Galois group
Recently, it has been proved in a precise and quantitative form that the monogenicity has an increasing effect on the class groups of number fields and orders; see Bhargava and Varma (2016), Ho, Shankar and Varma (2018), Bhargava, Hanke and Shankar (2020), Siad (2021), Swaminathan (2023).
The examples of degree $n=3,4,5,6$ in Section 6 show that the multiplicity of monogenicity can be relatively large if the Galois group is $S_{n}$, i.e. if its size is large relative to $n$.
Recently, Arpin, Bozlee, Herr and Smith (2023) studied so-called twisted monogenic relative extensions $K/L$. They proved that $L$ has trivial class group (this is the case if e.g. $L=\mathbb{Q}$) if and only if every twisted monogenic extension of $L$ is monogenic.
D. Distribution of monogenic number fields
As is well-known, all quadratic number fields are monogenic. For degree $n=3$, the first example of a non-monogenic number field was given by Dedekind (1878). For certain values $n\geq 3$, there are various results of the shape that there are infinitely many non-monogenic number fields of degree $n$.
Akhtari (2020) showed that a positive proportion of cubic number fields, when ordered by their absolute discriminant, are not monogenic. Alpöge, Bhargava and Shnidman (2020) proved more precisely that, if isomorphism classes of cubic number fields are ordered by their absolute discriminant, then a positive proportion are not monogenic and yet have no local obstruction to being monogenic. Recently, Alpöge, Bhargava and Shnidman (2024) proved a similar result for quartic number fields, and even that a positive proportion of quartic number fields are not rationally monogenic.
For $n=3,4,6$, tables of Gaál (2019) show that the frequency of monogenic number fields $K$ of degree $n$ is decreasing in tendency as the absolute value of the discriminant $|D_{K}|$ increases.
Denote by $N_{n}(X)$ the number of isomorphism classes of monogenic number fields $K$ of degree $n$ with $|D_{K}|\leq X$ and with associated Galois group $S_{n}$.
Theorem D.1 (Bhargava, Shankar and Wang, 2022).
For every $n\geq 2$ we have
$N_{n}(X)\gg X^{{1/2+1}/(n-1)}\ \ \text{as }X\to\infty.$
In the same paper, the authors conjecture that $X^{{1/2+1}/(n-1)}$ is the right order of magnitude.
E. Arithmetic characterization of monogenic and multiply monogenic number fields
The following problem continues to attract considerable attention:
Hasse’s problem (1960’s): give an arithmetic characterization of monogenic number fields.
In this direction there are many important results for deciding the monogenicity or non-monogenicity of number fields from certain special infinite classes, including cyclotomic, abelian, cyclic, pure, composite number fields, certain quartic, sextic, multiquadratic number fields and relative extensions, and parametric families of number fields defined by binomial, trinomial,… irreducible polynomials.
In their proofs various types of tools are used, among others Dedekind’s criterion; Newton polygons; Montes’ algorithm; Ore’s theorem; Engström’s theorem; Gröbner basis approach; reduction to binomial Thue equations; irreducible monic polynomials with square-free discriminant; non-squarefree discriminant approach; infinite parametric families of number fields; use of the index form approach.
For details, we refer to Dedekind (1878) and to the books Hensel (1908), Hasse (1963), Narkiewicz (1974), Evertse and Győry (2017), Gaál (2019) and the references given there. For some recent developments, see also the survey article Gaál (2024) with many interesting special results. We note that Hasse’s problem is not yet solved in full generality.
In general, while the monogenicity can be, its multiplicity cannot be determined by the arithmetic methods utilized.
Problem 5.
Give an arithmetic characterization of multiply monogenic number fields.
Clearly, Hasse’s problem and Problem 5 do not properly belong to the reduction theory of integral polynomials.
Dedekind’s necessary criterion for monogenicity of a number field was generalized by Del Corso, Dvornicich and Simon (2005) to a criterion for rational monogenicity. Perhaps this provides a tool to construct more examples of number fields that are not rationally monogenic.
Acknowledgment. We are very grateful to Dr. Csanád Bertok for typing a substantial part of the manuscript.